### v2

# Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping

Abstract
High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges. The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather. This paper proposes the PriorDrive framework to addresses these limitations by harnessing the power of prior maps, significantly enhancing the robustness and accuracy of online HD map construction. Our approach integrates a variety of prior maps, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data. To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements. At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs a dual encoding mechanism to process vector data. The intra-vector encoder captures fine-grained local features, while the inter-vector encoder integrates global context. Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance. Through extensive testing on the nuScenes dataset, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities. The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
Introduction
High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries (Elghazaly et al. 2023; Gao et al. 2020). Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive. These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments. In response, there is increasing interest in online HD map construction (Qiao et al. 2023; Ding et al. 2023; Zhang et al. 2024a), where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds. Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps. Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure LABEL:fig:priors. SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity. Existing HD maps (HD map-EX), despite their high accuracy, often fail to reflect current road conditions due to infrequent updates. Nevertheless, these maps offer valuable prior information for HD map construction. Additionally, historical prediction maps, which are often overlooked, can provide insights into current road conditions by reflecting the latest observations from previous perceptions. As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
Recent studies (Xiong et al. 2023; Jiang et al. 2024; Sun et al. 2023) have attempted to use prior maps to supplement perception data, but they face several limitations. For instance, NMP (Xiong et al. 2023) relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments. P-MapNet (Jiang et al. 2024) incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively. MapEX (Sun et al. 2023), on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps. These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model. However, designing such a model is challenging due to the variable length, sparsity, and discreteness of vector data. To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE). The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow the model to fully learn the prior distribution of vector maps. Our experiments on the nuScenes dataset demonstrate that our PriorDrive can be seamlessly applied to various online mapping models, resulting in significant performance improvements, including a +2.2 mIoU for HDMapNet and a +4.2 mAP for MapTRv2.
In summary, our contributions are as follows:
  • We introduce a unified vector encoder (UVE) that effectively encodes various vector data through a dual encoding mechanism: the Intra-Vector Encoder captures fine local features, while the Inter-Vector Encoder fuses global context information.
  • We propose a pre-training paradigm for vector data that learns the prior distribution of vector data by adding Gaussian noise or mask at the segment and point levels and reconstructing the entire vector map.
  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception. Our comprehensive evaluation on nuScenes dataset demonstrates that our PriorDrive significantly enhances the performance of online mapping models.
Related Work
Online Vectorized HD Map Construction
Recent advances in HD map construction have shifted from traditional offline methods based on SLAM techniques (Shan and Englot 2018; Shan et al. 2020) to more dynamic approaches that utilize vehicle-mounted sensors for online HD map generation (Xie, Pang, and Wang 2023; Chen, Deng, and Furukawa 2024; Liu et al. 2024a). This shift not only reduces costs but also ensures that maps reflect up-to-date road conditions. Some approaches (Li et al. 2021, 2022) treat HD map construction as a segmentation task, predicting pixel-level rasterized maps that require complex post-processing for vectorization. To address this limitation, recent work (Hu et al. 2024; Zhang et al. 2024b; Chen et al. 2024) has focused on developing end-to-end frameworks for vectorized map learning.
VectorMapNet (Liu et al. 2022) introduced a coarse-to-fine, two-stage network leveraging keypoint representations. MapTR (Liao et al. 2022) advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like (Carion et al. 2020) one-stage network. HIMap (Zhou et al. 2024) proposed a hybrid representation learning pipeline to capture both point-level and element-level information. MapQR (Liu et al. 2024c) introduced a scatter-and-gather query mechanism for precise detection of content and location information, while MGMap (Liu et al. 2024b) used mask-guided features to refine lane shape representations with enhanced feature detail. However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
Online Mapping Based on Prior Maps
Recent research has explored the integration of prior maps to enhance the performance of online mapping models. P-MapNet (Jiang et al. 2024) encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps. NMP (Xiong et al. 2023) introduces a global neural map prior that self-updates, thereby improving the performance of local map inference. MapEX (Sun et al. 2023) categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm. Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information. Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
Pretrained Methods Based on Mask Modeling
In the fields of Natural Language Processing and Computer Vision, masked pre-training has proven to be an effective strategy for self-supervised representation learning. In NLP, models like BERT (Devlin et al. 2019; Liu et al. 2019; Lan et al. 2019) use masked language modeling to predict randomly masked tokens within a bidirectional text context. Similarly, methods such as MAE (He et al. 2021; Li et al. 2023; Liu et al. 2023) mask random patches of input images and reconstruct them based on the remaining unmasked patches. Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data. Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized map data.
Proposed Method: PriorDrive
Figure 1: Our Integration Method. For query-based models, we enable interaction between the prior features and queries at both the instance-level and point-level, ensuring a more nuanced integration. Hierarchical query comes from MapTRv2 (Liao et al. 2023). For the other models, we integrate the prior features directly with the BEV features.
Problem Formulation
Formulation of Online Mapping.
The goal of online mapping is to construct a local HD map using on-board sensor observations, such as the surrounding camera images. The online mapping process typically involves two main modules: the Bird’s-Eye View (BEV) encoder and the map decoder. The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features. The map decoder usually comprises a Fully Convolutional Network (FCN) (Shelhamer, Long, and Darrell 2015) or a query-based module similar to DETR (Carion et al. 2020), which outputs either rasterized perception results or vectorized map elements.
Let the set of images be denoted as $\{I_{1},\ldots,I_{k}\}$, and the BEV encoder be represented as ${E}_{bev}$. The extraction of BEV features ${f}_{bev}$ can then be expressed as:
$f_{bev}={E}_{bev}(\{I_{1},\ldots,I_{K}\}).$ (1)
Similarly, let the map decoder module be denoted as $D_{map}$. The final process of predicting map elements can be expressed as:
$\mathcal{P}={D}_{map}(f_{bev},~{}opt({Q})),$ (2)
where $\mathcal{P}$ represents the predicted map elements, $Q$ denotes the learnable query, and $opt$ indicates that this learnable query is optional, which is required in query-based models but not in other models.
Extraction of Prior Map Features.
Our proposed UVE serves as a unified vector encoder that can directly encode various vector data information, such as the position ($x$, $y$), direction ($v_{x}$, $v_{y}$), and type of points ($c$). Using $p=\left[x,y,v_{x},v_{y},c\right]$ to represent the point, a vector $v$ typically consists of multiple ordered sets of points, $v=\left[p_{1},p_{2},\ldots,p_{n}\right]$, where $n$ varies. The map prior $M_{prior}$ consists of multiple vectors $M_{prior}=\{v_{1},v_{2},\ldots,v_{m}\}$, with $m$ being variable. Using $E_{uve}$ to represent the UVE model, the extraction process of the features of the prior map, denoted as $f_{prior}$, can be represented as:
$f_{prior}=E_{uve}(M_{prior}).$ (3)
Online Mapping Based on Prior Maps.
We proposed four methods in total to integrate prior map features into online mapping models. For the methods without a learnable query $Q$, such as HDMapNet (Li et al. 2021), we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{BEV}$. Then connect it with the BEV feature $f_{BEV}$ and aligned by $Conv$, as shown in Figure 1 left.
$f_{bev}^{*}=Conv(cat[f_{bev},~{}Deconv(reshape(f_{prior}))]),$ (4)
here, $f_{bev}^{*}$ represents the BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
For methods based on learnable query $Q$, such as the vectorized model MapTR series (Liao et al. 2023), we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 1 right. Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively. The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$. So we interact prior features and queries at the instance level and click, respectively. The HPQuery of addition, replacement, and concatenation operations can be expressed as:
$q_{ij}^{*}=q_{i}^{ins}.add(f_{i}^{ins})+q_{j}^{pt}.add(f_{j}^{pt}),$ (5)
$q_{ij}^{*}=q_{i}^{ins}.replace(f_{i}^{ins})+q_{j}^{pt}.replace(f_{j}^{pt}),$ (6)
$q_{ij}^{*}=concat[q_{i}^{ins},~{}f_{i}^{ins}]+concat[q_{j}^{pt},~{}f_{j}^{pt}].$ (7)
We will detail the effectiveness of these feature fusion methods in the experimental section.
Figure 2: Structure of UVE. Hybrid prior embedding enhances vector representation. Intra-instance and inter-instance attention mechanisms are employed to refine local features and capture global context effectively.
Architecture of UVE
Inspired by BERT (Devlin et al. 2019) on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 2). And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
Hybrid Prior Embedding.
As shown in Figure 2, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE). We use the method from (Tancik et al. 2020) for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$. We concatenate these two embeddings as the point-level embedding. To obtain the feature representation of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding. Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances. To ensure the order of points, we also introduce 2D learnable position embedding. Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
Intra and Inter-Vector Encoder.
Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector. As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances. Through this dual encoding mechanism, UVE can deeply understand and represent vector data, providing strong support for key tasks such as online HD maps construction.
Figure 3: Pre-trained Pipeline. The noise & mask generator creates segment-level and point-level noise or mask, and then reconstructs the entire map via UVE and MLP.
Pre-training UVE: Position Modeling
Due to the limited inference ability of online mapping model, there are errors in historical prediction maps. So we used position modeling to pre-train UVE to improve its encoding and noise reduction capabilities. specifically, there are two methods: random noise and masking (see Figure 3).
Noise & Mask Generator.
The noise is primarily categorized into segment-level and point-level. In point-level noise, random noise is added to $5\%$ of the vector points across the entire map. For segment-level noise, random noise is added to all vector points within a sub-vector segment, after randomly selecting $10\%$ of the map elements. Using $M_{org}$ to represent the original map elements (with the same data structure as $M_{prior}$), the process of adding Gaussian noise can be represented by the following formula:
$M_{org}^{*}=M_{org}\emph{[random\_index]}+\epsilon,$ (8)
where $\epsilon\sim\mathcal{N}(0,1)$ is a random noise, which is only added to the horizontal and vertical coordinates of each point.
Similar to adding noise to the points, the selection of points for adding the mask is also divided into point-level and segment-level. After selecting the points to add the mask, the coordinates of these points will be masked as $-1$.
[…]
Loss Function.
We feed the vector map to UVE encoding after passing through the noise mask generator, then decode the coordinates of all vector points using MLP, and calculate mean euclide distance from real coordinates as supervision:
$\mathcal{L}=RMSE(P,~{}mlp(E_{uve}(M_{org}^{*}))),$ (10)
where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in the $M_{org}$.
Experiments
Experimental Settings
Benchmark.
We mainly validate our method on the popular nuScenes (Caesar et al. 2019) dataset following the standard setting of previous methods (Li et al. 2021; Liu et al. 2022; Liao et al. 2022). NuScenes is a large-scale autonomous driving benchmark, that includes multiple traversals with precise localization and annotated HD maps labels. It contains a total of 1,000 scenes, each with an approximate duration of 20 seconds. The data was collected utilizing a 32-beam LiDAR operating at 20Hz and six cameras offering a 360-degree field of view at 12 Hz. Annotations for keyframes are furnished at a rate of 2Hz, with the training and validation sets containing 28,130 frames and 6,019 frames, respectively.
Evaluation Protocol.
For the vectorized results, we follow the standard metric used in previous works (Li et al. 2021; Liu et al. 2022; Liao et al. 2022). The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis. And we adopt average precision (AP) evaluated the quality of map construction. Chamfer distance $D_{Chamfer}$ is utilized to determine whether the prediction and ground truth (GT) are matched or not. We compute the $\mathrm{AP}_{\tau}$ under several $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
$\mathrm{AP}=\frac{1}{|\mathrm{T}|}\sum_{\tau\in\mathrm{T}}\mathrm{AP}_{\tau}.$ (11)
Following the previous methods (Li et al. 2021; Liu et al. 2022; Liao et al. 2022), we have selected three kinds of map elements are chosen for fair evaluation: lane divider, pedestrian crossing and road boundary. Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works (Li et al. 2021; Xiong et al. 2023; Jiang et al. 2024) and employ intersection-over-union (IoU) as the metrics for segmentation results.
Implementation Details.
We primarily perform our experiments using the latest and influential vectorized model MapTRv2 (Liao et al. 2023). In order to validate the broad applicability of our method, we also use HDMapNet (Li et al. 2021) as baseline. For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes. All experiments were trained with 6 NVIDIA RTX A6000. For pre-training UVE, we trained the 24-epoch on nuScenes vecror data. For historical prediction maps, we run each base model to get their respective online local maps.
Method Modality Backbone Epoch Metric ped. div. bou. mean $\mathbf{\tau_{0.5m}}$
HDMapNet C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
+ SD map Prior C Effi-B0 30 IoU 20.8 41.7 41.8 34.8 -
+ Online Local Map Prior C Effi-B0 30 IoU 22.5 41.3 41.7 35.1 -
+ HD map-EX* Prior C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
MapTRv2 C R50 24 AP 59.8 62.4 62.4 61.5 38.6
+ SD map Prior C R50 24 AP 61.7 65.5 66.4 64.5 42.7
+ Online Local Map Prior C R50 24 AP 63.6 67.3 66.2 65.7 44.7
+ HD map-EX* Prior C R50 24 AP 61.5 65.6 68.3 65.1 44.3
Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range. “C” denotes camera input. The best results using the same backbone are highlighted in bold. The SD map is sourced from OpenStreetMap (Haklay and Weber 2008). The online local map refers to the historical prediction results. As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX (Sun et al. 2023) to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
M N Dim IoU
Intra Inter ped. div. bou. mean
4 4 64 20.7 41.3 41.1 34.4
6 2 64 20.5 41.1 40.9 34.2
2 6 64 20.7 41.0 41.3 34.3
2 2 128 19.5 41.4 41.2 34.0
2 2 64 20.8 41.7 41.8 34.8
1 1 32 19.8 41.1 40.9 33.9
Table 2: Ablation experiments on UVE structure within HDMapNet and SD map prior.
Method Dist Err. AP
Type Seg Pt ped. div. bou. mean
Without Pretrain - 61.3 63.4 64.4 63.0
Original - - 0.09m 61.2 64.1 64.5 63.3
Noise 3m 20% - 0.46m 61.4 63.0 63.8 62.7
Noise 1m 10% - 0.15m 61.6 63.8 64.2 63.2
Noise 1m 10% 5% 0.18m 61.6 64.3 64.8 63.6
Mask 10% 5% 0.43m 61.8 64.0 64.3 63.4
Table 3: Ablation experiments of the pre-training strategies on MapTRv2. Here, we fixed prior_num=2, search_range=5m and integration=replace. “Seg” and “Pt” represent segment-level and point-level, respectively.
Main Results
To evaluate the effectiveness of our method across different model architectures, evaluation metrics, and vector prior maps, we integrated the UVE-encoded prior maps into both HDMapNet and the vectorized model MapTRv2 (see Table 1). The prior information provided by SD maps, which primarily consists of a centerline skeleton, resulted in the smallest improvement. Specifically, MapTRv2 and HDMapNet showed enhancements of 3.0 mAP and 1.9 mIoU, respectively, when using SD map priors. And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior information. The incorporation of online local maps, which offer up-to-date information, led to more significant improvements. This is reflected in the enhancements of 4.2 mAP and 2.2 mIoU for MapTRv2 and HDMapNet, respectively. These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information. Figure 4 illustrates how online local maps effectively restore obscured lane dividers, pedestrian crossings, and road boundaries. Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively. Additionally, we evaluated the results using a stricter threshold of 0.5m. Under this criterion, our method with SD map, online local map, and outdated HD map (HD map-EX*) priors achieved improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively. These results underscore that our method provides more substantial improvements under stricter evaluation conditions. Overall, our method demonstrated consistent performance improvements across all baseline models with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
Ablation Study
In this section, we do extensive research on UVE structure, pre-training, number and search range of prior maps, and integration methods. we conduct experiments using online local map priors on MapTRv2 unless otherwise specified.
Structure of UVE.
Table 2 presents the results of ablation experiments on the UVE structure using the SD prior map within HDMapNet. We observed that increasing the number of layers or expanding the feature dimension of UVE yielded only marginal improvements, averaging around 1.3 mIoU. This limited gain may be attributed to the model’s excess capacity, leading to overfitting. Conversely, a minimal configuration of UVE, with only one layer each of intra-vector and inter-vector attention and a dimension of 32, resulted in a modest improvement of just 1 mIoU. However, when UVE was configured with two layers of intra-vector and inter-vector attention and a feature dimension of 64, the model’s performance reached saturation.
The Effect of Pre-training.
To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies, as shown in Table 3. When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters. However, the simplicity of the task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement. We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly. The UVE’s average Euclidean distance error in reconstructing the entire map was 0.46 meters. However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged. Next, we applied 1-meter segment-level noise to 10% of the map elements randomly, resulting in a reduced error of 0.15 meters for UVE and a 0.2 mAP improvement compared to the model without pre-training. Building on this, we introduced 1-meter point-level noise to 5% of the vector points. Although this increased UVE’s error to 0.18 meters, it yielded a more substantial mAP improvement of 0.6. This demonstrates that pre-training strategies incorporating both segment-level and point-level noise are more effective at learning the prior information of vector maps. We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement. This suggests that while challenging tasks based on masking can pre-train UVE effectively, their impact may be constrained by the amount of pre-training data.
Impact of the Number of Prior Maps.
We conducted an ablation study to assess the effect of varying the number of prior maps on model performance (see Table 4). The results indicate that using a single prior map led to a 0.9 mAP improvement, demonstrating its effectiveness in enhancing current perception. Incorporating two prior maps resulted in a more substantial improvement of 1.5 mAP, reflecting the value of richer prior information. However, using multiple prior maps only yielded a 1.3 mAP improvement. This could be attributed to the integration method, where replacing too many original queries with prior features may have diminished the performance of the original queries.
Effect of Different Search Ranges for Prior Maps.
We performed ablation experiments to explore the impact of search ranges for prior maps in Table 5. The results indicate that increasing the search range leads to a slight decrease in improvement. This suggests that while a larger search range allows the prior map to cover a wider area, it also introduces greater positional error. On the other hand, a more precise, narrower search range may not always provide multiple prior maps for most locations, limiting the benefits.
Different Methods for Integrating Prior Maps.
To evaluate the versatility of our approach, we tested various integration methods in Table 6. Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.1 and 1.6, respectively, but these methods may impact the performance of the original query. In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP. This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
Inference Speed, Memory, Model Size, and Detailed Runtime.
Table 7 compares our method with MapTRv2 in the above aspects. The results show that our method maintains similar FPS, GPU memory usage, and parameter count compared to MapTRv2, while achieving a 4.2 mAP improvement. Additionally, we provide the detailed runtime of each component. Our proposed UVE accounts for only 15.3% of the total pipeline runtime, with the majority of inference time spent on the 2D Backbone (50.2%), highlighting the efficiency and practical value of our approach across various application scenarios.
Fixed Prior AP
Parameters Num ped. div. bou. mean
- None 59.8 62.4 62.4 61.5
search_range=5m One 61.9 62.1 62.4 62.4
pretrain=false Two 61.3 63.4 64.4 63.0
integration=replace Multiple 61.2 63.0 64.2 62.8
Table 4: Ablation experiments on the number of online local prior maps on MapTRv2.
Fixed Search AP
Parameters Range ped. div. bou. mean
- None 59.8 62.4 62.4 61.5
prior_num=2 5 m 61.3 63.4 64.4 63.0
pretrain=false 10 m 60.8 63.7 64.0 62.8
integration=replace 15 m 61.5 62.7 64.0 62.7
Table 5: Ablation experiments on the search range of online local prior maps on MapTRv2.
Fixed Integrated AP
Parameters Method ped. div. bou. mean
- None 59.8 62.4 62.4 61.5
prior_num=2 Replace 61.6 64.3 64.8 63.6
pretrain=true Add 62.5 63.3 63.6 63.1
search_range=5m Concat 63.6 67.3 66.2 65.7
Table 6: Ablation experiments on the integrated methods of online local prior maps on MapTRv2. Note that we proposed other integrated methods after pre-training UVE, so we use pre-trained UVE here.
- MapTRv2 Ours
FPS 10.3 9.9
GPU mem. (MB) 2624 2650
Params (MB) 40.3 43.4
mAP 61.5 65.7
Component Runtime (ms) Proportion
2D Backbone 47.6 50.2%
PV to BEV Module 12.9 13.6%
UVE 14.6 15.3%
Map Decoder 19.8 20.9%
Table 7: Detailed comparison of our method with MapTRv2. FPS and detailed runtime of each component are measured on one A6000 GPU with batch size as 1.
Figure 4: Qualitative results with and w/o online local prior. Prior maps can help restore obscured map elements and make predictions more complete and accurate.
Conclusion
In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction for autonomous vehicles. Central to our approach is the UVE, which we designed to encode diverse vector data efficiently. Through comprehensive experimentation, we demonstrated that UVE, combined with our proposed pre-training strategies, significantly improves the performance of state-of-the-art mapping models. Our approach not only addresses the challenges associated with real-time HD map construction in dynamic and complex environments but also offers a scalable solution that continually improves map accuracy over time. The iterative use of historical prediction maps as priors leads to progressively refined map outputs, thus ensuring that the most up-to-date road information is available for autonomous navigation.

### v3

@@
 To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements.
-At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs a dual encoding mechanism to process vector data.
-The intra-vector encoder captures fine-grained local features, while the inter-vector encoder integrates global context.
+At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs hybrid prior embedding and a dual encoding mechanism to process vector data.
 Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance.
-Through extensive testing on the nuScenes dataset, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities.
+Through extensive testing on the nuScenes, Argoverse 2 and OpenLane-V2, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities.
 The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
-Introduction
-High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries (Elghazaly et al.
-2023; Gao et al.
-2020).
+Figure 1: Overview of PriorDrive.
+Our unified vector encoder (UVE) directly encodes diverse vector prior maps and seamlessly integrates them into existing online mapping frameworks.
+This integration enhances the final predictions, making them more complete and accurate than those generated without prior information.
+footnotetext: Work done during the internship at Amap, Alibaba Group.footnotetext: Equal contribution.footnotetext: Corresponding author.
+1 Introduction
+High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries [6, 8].
 Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive.
 These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments.
-In response, there is increasing interest in online HD map construction (Qiao et al.
-2023; Ding et al.
-2023; Zhang et al.
-2024a), where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds.
+In response, there is increasing interest in online HD map construction [27, 5, 38], where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds.
 Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
 In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps.
-Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure LABEL:fig:priors.
+Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure 1.
 SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity.
@@
 As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
-Recent studies (Xiong et al.
-2023; Jiang et al.
-2024; Sun et al.
-2023) have attempted to use prior maps to supplement perception data, but they face several limitations.
-For instance, NMP (Xiong et al.
-2023) relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments.
-P-MapNet (Jiang et al.
-2024) incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively.
-MapEX (Sun et al.
-2023), on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps.
+Recent studies [35, 13, 29] have attempted to use prior maps to supplement perception data, but they face several limitations.
+For instance, NMP [35] relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments.
+P-MapNet [13] incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively.
+MapEX [29], on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps.
 These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
+In addition to constructing accurate online HD maps, the autonomous driving scenario requires understanding the topology in road scenes (i.e., lane connectivity and correspondence with traffic elements), which is called topology reasoning task.
+Since connected SD maps include crucial information for road topology, we also apply PriorDrive to this task to enhance the performance of topology reasoning.
 Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model.
@@
 To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE).
-The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow the model to fully learn the prior distribution of vector maps.
-Our experiments on the nuScenes dataset demonstrate that our PriorDrive can be seamlessly applied to various online mapping models, resulting in significant performance improvements, including a +2.2 mIoU for HDMapNet and a +4.2 mAP for MapTRv2.
+The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow model to fully learn the prior distribution of vector maps.
+Our experiments on nuScenes, Argoverse 2 and OpenLane-V2 demonstrate that our PriorDrive is plug-and-play and can be seamlessly applied to various online mapping models and vector tasks, resulting in significant performance improvements.
 In summary, our contributions are as follows:
-  • We introduce a unified vector encoder (UVE) that effectively encodes various vector data through a dual encoding mechanism: the Intra-Vector Encoder captures fine local features, while the Inter-Vector Encoder fuses global context information.
-  • We propose a pre-training paradigm for vector data that learns the prior distribution of vector data by adding Gaussian noise or mask at the segment and point levels and reconstructing the entire vector map.
+  • We introduce a unified vector encoder that effectively captures diverse vector attributes and enables encoding of variable-length vectors into fixed-length features through Hybrid prior embedding and dual encoding mechanism.
+  • We first proposed a pre-training paradigm for vector data, which can effectively learn the prior distribution of vector data and help reduce noise in historical maps through point-level and segment-level pre-training.
   • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception.
-Our comprehensive evaluation on nuScenes dataset demonstrates that our PriorDrive significantly enhances the performance of online mapping models.
-Related Work
-Online Vectorized HD Map Construction
-Recent advances in HD map construction have shifted from traditional offline methods based on SLAM techniques (Shan and Englot 2018; Shan et al.
-2020) to more dynamic approaches that utilize vehicle-mounted sensors for online HD map generation (Xie, Pang, and Wang 2023; Chen, Deng, and Furukawa 2024; Liu et al.
-2024a).
-This shift not only reduces costs but also ensures that maps reflect up-to-date road conditions.
-Some approaches (Li et al.
-2021, 2022) treat HD map construction as a segmentation task, predicting pixel-level rasterized maps that require complex post-processing for vectorization.
-To address this limitation, recent work (Hu et al.
-2024; Zhang et al.
-2024b; Chen et al.
-2024) has focused on developing end-to-end frameworks for vectorized map learning.
-VectorMapNet (Liu et al.
-2022) introduced a coarse-to-fine, two-stage network leveraging keypoint representations.
-MapTR (Liao et al.
-2022) advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like (Carion et al.
-2020) one-stage network.
-HIMap (Zhou et al.
-2024) proposed a hybrid representation learning pipeline to capture both point-level and element-level information.
-MapQR (Liu et al.
-2024c) introduced a scatter-and-gather query mechanism for precise detection of content and location information, while MGMap (Liu et al.
-2024b) used mask-guided features to refine lane shape representations with enhanced feature detail.
+Our comprehensive evaluation demonstrates that PriorDrive is plug-and-play and significantly enhances the performance of online mapping models.
+2 Related work
+2.1 Online vectorized HD map construction
+Online vectorized HD map construction was initially treated as a segmentation task [15, 18].
+To construct vectorized HD map, HDMapNet [15] requires complex post-processing of pixel-level rasterized maps.
+VectorMapNet [25] introduced a coarse-to-fine, two-stage network leveraging keypoint representations.
+MapTR [19] advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like [2] one-stage network.
+Subsequent studies [41, 39, 36, 9, 12, 40, 3] have introduced various insightful approaches for further improvement.
 However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
-Online Mapping Based on Prior Maps
+The release of OpenLane-V2 [31] dataset has sparked growing interest in topology reasoning within driving scenes.
+This task focuses on recognizing the topologies among lanes and lanes with traffic elements.
+TopoNet [17] leverages GNN to update lane representation and lane topology.
+TopoMLP [34] utilizes PETR [24] for centerline detection and employs simple MLPs to predict the relationships.
+In this paper, we also apply PriorDrive to topological inference tasks, introducing interconnected SD Map as an additional input to support topology reasoning.
+2.2 Online mapping based on prior maps
 Recent research has explored the integration of prior maps to enhance the performance of online mapping models.
-P-MapNet (Jiang et al.
-2024) encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps.
-NMP (Xiong et al.
-2023) introduces a global neural map prior that self-updates, thereby improving the performance of local map inference.
-MapEX (Sun et al.
-2023) categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm.
+P-MapNet [13] encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps.
+NMP [35] introduces a global neural map prior that self-updates, thereby improving the performance of local map inference.
+MapEX [29] categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm.
 Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information.
 Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
-Pretrained Methods Based on Mask Modeling
-In the fields of Natural Language Processing and Computer Vision, masked pre-training has proven to be an effective strategy for self-supervised representation learning.
-In NLP, models like BERT (Devlin et al.
-2019; Liu et al.
-2019; Lan et al.
-2019) use masked language modeling to predict randomly masked tokens within a bidirectional text context.
-Similarly, methods such as MAE (He et al.
-2021; Li et al.
-2023; Liu et al.
-2023) mask random patches of input images and reconstruct them based on the remaining unmasked patches.
+2.3 Pretrained methods based on mask modeling
+In the fields of NLP and CV, masked pre-training has proven to be an effective strategy for self-supervised representation learning.
+In NLP, models like BERT [4, 23, 14] use masked language modeling to predict randomly masked tokens within a bidirectional text context.
+Similarly, methods such as MAE [11, 16, 21] mask random patches of input images and reconstruct them based on the remaining unmasked patches.
 Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data.
-Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized map data.
-Proposed Method: PriorDrive
+Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized data.
+3 Proposed method: PriorDrive
 Figure 2: Our Integration Method.
-For query-based models, we enable interaction between the prior features and queries at both the instance-level and point-level, ensuring a more nuanced integration.
-Hierarchical query comes from MapTRv2 (Liao et al.
-2023).
-For the other models, we integrate the prior features directly with the BEV features.
-Problem Formulation
+For query-based models, we enable interaction between the prior features and queries at both instance-level and point-level, ensuring a more nuanced integration.
+Hierarchical query comes from MapTRv2 [20].
+For the other models, we integrate prior features directly with the BEV features.
+3.1 Problem formulation
 Formulation of online mapping.
@@
 The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features.
-The map decoder usually comprises a Fully Convolutional Network (FCN) (Shelhamer, Long, and Darrell 2015) or a query-based module similar to DETR (Carion et al.
-2020), which outputs either rasterized perception results or vectorized map elements.
+The map decoder usually comprises a Fully Convolutional Network [28] or a query-based module similar to DETR [2], which outputs either rasterized perception results or vectorized map elements.
 Let the set of images be denoted as $\{I_{1},\ldots,I_{k}\}$, and the BEV encoder be represented as ${E}_{bev}$. The extraction of BEV features ${f}_{bev}$ can then be expressed as:
@@
 We proposed four methods in total to integrate prior map features into online mapping models.
-For the methods without a learnable query $Q$, such as HDMapNet (Li et al.
-2021), we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{BEV}$.
+For the methods without a learnable query $Q$, such as HDMapNet [15], we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{bev}$.
 Then connect it with the BEV feature $f_{bev}$ and aligned by $Conv$, as shown in Figure 2 left.
 $f_{bev}^{*}=Conv(cat[f_{bev},~{}Deconv(reshape(f_{prior}))]),$ (4)
-here, $f_{bev}^{*}$ represents the BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
-For methods based on learnable query $Q$, such as the vectorized model MapTR series (Liao et al.
-2023), we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 1 right.
+here, $f_{bev}^{*}$ represents BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
+For methods based on learnable query $Q$, such as the vectorized model MapTR series [20], we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 2 right.
 Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively.
@@
 Intra-instance and inter-instance attention mechanisms are employed to refine local features and capture global context effectively.
-Architecture of UVE
-Inspired by BERT (Devlin et al.
-2019) on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 2).
+3.2 Architecture of UVE
+Inspired by BERT [4] on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 3).
 And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
@@
 As shown in Figure 3, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE).
-We use the method from (Tancik et al.
-2020) for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$.
+We use the method from [30] for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$.
 We concatenate these two embeddings as the point-level embedding.
@@
 Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
-Intra and Inter-Vector Encoder.
+Dual encoding mechanism.
 Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector.
 As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances.
-Through this dual encoding mechanism, UVE can deeply understand and represent vector data, providing strong support for key tasks such as online HD maps construction.
 Figure 4: Pre-trained Pipeline. The noise & mask generator creates segment-level and point-level noise or mask, and then reconstructs the entire map via UVE and MLP.
-Pre-training UVE: Position Modeling
+3.3 Pre-training UVE: Position modeling
 Due to the limited inference ability of online mapping model, there are errors in historical prediction maps.
@@
 $\mathcal{L}=RMSE(P,~{}mlp(E_{uve}(M_{org}^{*}))),$ (10)
-where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in the $M_{org}$.
-Experiments
-Experimental Settings
+where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in $M_{org}$.
+4 Experiments
+4.1 Experimental settings
+Method Modality Backbone Epoch Metric ped. div. bou. mean $\mathbf{\tau_{0.5m}}$
+HDMapNet [15] C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
+# SD map Prior (Ours) C Effi-B0 30 IoU 22.2 42.0 41.5 35.2 -
+# Online Local Map Prior (Ours) C Effi-B0 30 IoU 23.1 41.4 42.0 35.5 -
+# HD map-EX* Prior (Ours) C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
+MapTR [19] C R50 24 AP 41.2 49.5 51.1 47.3 -
+# P-MapNet [13] C R50 24 AP 43.7 50.9 53.5 49.4 -
+StreamMapNet [37] C R50 24 AP 60.4 61.9 58.9 60.4 -
+# SQD-MapNet [32] C R50 24 AP 63.0 65.5 63.3 63.9 -
+MGMap [22] C R50 24 AP 61.8 65.0 67.5 64.8 -
+MapTRv2 [20] C R50 24 AP 59.8 62.4 62.4 61.5 38.6
+# SD map Prior (Ours) C R50 24 AP 61.7 65.5 66.4 64.5 42.7
+# Online Local Map Prior (Ours) C R50 24 AP 63.6 67.3 66.2 65.7 44.7
+# HD map-EX* Prior (Ours) C R50 24 AP 61.5 65.6 68.3 65.1 44.3
+Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range.
+“C” denotes camera input.
+The SD map is sourced from OpenStreetMap [10].
+The online local map refers to the historical prediction results.
+As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX [29] to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
 Benchmark.
-We mainly validate our method on the popular nuScenes (Caesar et al.
-2019) dataset following the standard setting of previous methods (Li et al.
-2021; Liu et al.
-2022; Liao et al.
-2022).
-NuScenes is a large-scale autonomous driving benchmark, that includes multiple traversals with precise localization and annotated HD maps labels.
-It contains a total of 1,000 scenes, each with an approximate duration of 20 seconds.
-The data was collected utilizing a 32-beam LiDAR operating at 20Hz and six cameras offering a 360-degree field of view at 12 Hz.
-Annotations for keyframes are furnished at a rate of 2Hz, with the training and validation sets containing 28,130 frames and 6,019 frames, respectively.
+We mainly validate our method on the popular nuScenes [1] and Argoverse 2 [33] following the standard setting of previous methods [25, 19].
+NuScenes contains a total of 1,000 scenes, each with an approximate duration of 20 seconds.
+The data was collected utilizing a 32-beam LiDAR operating and six cameras offering a 360-degree field of view.
+Argoverse 2 includes 1000 logs, each offering 15 seconds of 20Hz RGB images captured from 7 cameras, along with a 3D vector map.
+Furthermore, to demonstrate that PriorDrive is plug-and-play and widely applicable to other vector tasks, we conducted a topology inference task on OpenLane-V2 [31] dataset.
+Subset A, developed by Argogorse 2 [33], includes seven views as input, along with an additional SD map.
 Evaluation protocol.
-For the vectorized results, we follow the standard metric used in previous works (Li et al.
-2021; Liu et al.
-2022; Liao et al.
-2022).
+For the vectorized results, we follow the standard metric used in previous works [15, 25, 19].
 The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis.
 And we adopt average precision (AP) evaluated the quality of map construction.
-Chamfer distance $D_{Chamfer}$ is utilized to determine whether the prediction and ground truth (GT) are matched or not.
-We compute the $\mathrm{AP}_{\tau}$ under several $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
+We compute the $\mathrm{AP}_{\tau}$ under several Chamfer distance $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
 $\mathrm{AP}=\frac{1}{|\mathrm{T}|}\sum_{\tau\in\mathrm{T}}\mathrm{AP}_{\tau}.$ (11)
-Following the previous methods (Li et al.
-2021; Liu et al.
-2022; Liao et al.
-2022), we have selected three kinds of map elements are chosen for fair evaluation: lane divider, pedestrian crossing and road boundary.
-Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works (Li et al.
-2021; Xiong et al.
-2023; Jiang et al.
-2024) and employ intersection-over-union (IoU) as the metrics for segmentation results.
+Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works [15, 35, 13] and employ IoU as the metrics for segmentation results.
+For OpenLane-V2’s lane centerline perception task, $DET_{l}$ averages the Frechet distance to quantify similarity.
+$DET_{t}$ uses IoU and averages it across various traffic categories.
+$TOP_{ll}$ and $TOP_{lt}$ compute the topology matrix similarity between lanes and between lanes and traffic elements, respectively.
+The overall metric is called OLS:
+$\mathrm{OLS}=\frac{1}{4}[\mathrm{DET}_{l}+\mathrm{DET}_{t}+f(\mathrm{TOP}_{ll})+f(\mathrm{TOP}_{lt})],$ (12)
+where $f$ is the square root function.
 Implementation details.
-We primarily perform our experiments using the latest and influential vectorized model MapTRv2 (Liao et al.
-2023).
-In order to validate the broad applicability of our method, we also use HDMapNet (Li et al.
-2021) as baseline.
 For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes.
@@
 For historical prediction maps, we run each base model to get their respective online local maps.
-Method Modality Backbone Epoch Metric ped. div. bou. mean $\mathbf{\tau_{0.5m}}$
-HDMapNet C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
-+ SD map Prior C Effi-B0 30 IoU 20.8 41.7 41.8 34.8 -
-+ Online Local Map Prior C Effi-B0 30 IoU 22.5 41.3 41.7 35.1 -
-+ HD map-EX* Prior C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
-MapTRv2 C R50 24 AP 59.8 62.4 62.4 61.5 38.6
-+ SD map Prior C R50 24 AP 61.7 65.5 66.4 64.5 42.7
-+ Online Local Map Prior C R50 24 AP 63.6 67.3 66.2 65.7 44.7
-+ HD map-EX* Prior C R50 24 AP 61.5 65.6 68.3 65.1 44.3
-Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range.
-“C” denotes camera input.
-The best results using the same backbone are highlighted in bold.
-The SD map is sourced from OpenStreetMap (Haklay and Weber 2008).
-The online local map refers to the historical prediction results.
-As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX (Sun et al.
-2023) to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
+Method $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
+MGMap [22] 52.8 67.5 68.1 62.8
+MapQR [26] 60.1 71.2 66.2 65.9
+MapTRv2 [20] 60.7 68.9 64.5 64.7
+# SD Prior (Ours) 62.0 70.2 66.3 66.2
+# Online Local Prior (Ours) 63.9 71.7 68.6 68.1
+# HD map-EX* (Ours) 61.5 69.8 69.4 66.9
+Table 2: Comparison on Argoverse 2 where the 3D map elements are predicted directly.
+4.2 Main results
+Results on nuScenes.
+To evaluate the effectiveness of our method across different model, evaluation metrics, and vector prior maps, we choose HDMapNet and MapTRv2 as baselines (see Table 1).
+SD maps, which primarily consists of a centerline skeleton, showed enhancements of 3.0 mAP for MapTRv2 and 2.3 mIoU for HDMapNet, respectively.
+And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior maps.
+Online local maps, which offer up-to-date information, led to more significant improvements, with 4.2 mAP and 2.6 mIoU for MapTRv2 and HDMapNet, respectively.
+These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information.
+Figure 5 illustrates how online local maps effectively restore obscured map elements.
+More visualization are shown in the supplementary material.
+Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively.
+Additionally, we evaluated the results using a stricter threshold of 0.5m, achieving improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively, underscoring that our method provides more substantial improvements under stricter evaluation conditions.
+Overall, our method demonstrated consistent performance improvements across all baselines with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
+Method $\textbf{DET}_{l}$ $\textbf{DET}_{t}$ $\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS
+MapTR [19] 17.7 43.5 5.9 15.1 31.0
+TopoNet [17] 28.6 48.6 10.9 23.8 39.8
+TopoLogic [7] 29.9 47.2 23.9 25.4 44.1
+# SD Prior (Ours) 29.9 47.1 29.6 32.9 47.2
+# Online Local Prior (Ours) 31.0 47.8 30.9 32.5 47.9
+Table 3: Performance comparison with SOTA methods on OpenLaneV2 subset_A set, where SD Prior comes from the SD map of the openLanev2 itself.
+Since the map elements here have only centerline and no boundaries, there is no HD map-EX* prior.
+Results on Argoverse 2.
+In Table 2, we present the results of Argoverse 2, which provides 3D vectorized map elements with additional height information as ground truth.
+These testings are all trained 6 epochs with ResNet50 as the backbone.
+Our PriorDrive improves performance by 1.5, 3.4, and 2.2, respectively, using different prior maps, demonstrating the superior generalization ability of PriorDrive in terms of 3D vectorized map construction.
+Results on OpenLane-V2.
+We compared PriorDrive with SOTA methods when using ResNet-50 for training 24 epochs.
+Table 3 shows the results on subset A of OpenLane-V2.
+Compared to TopoLogic, our method achieves better topological reasoning accuracy using connected SD priors (29.6 v.s.
+23.9 on $TOP_{ll}$, 32.9 v.s.
+25.4 on $TOP_{lt}$), and the overall score OLS increased by 3.1, indicating that connected SD is more beneficial for lane connectivity and correspondence with traffic elements.
+Employing online local priors resulted in decent detection accuracy for both detection and topological reasoning, with OLS rising to 47.9.
+Overall, PriorDrive is a plug-and-play solution that can be widely applied to other vector tasks.
 M N Dim IoU
@@
 Table 4: Ablation experiments on UVE structure within HDMapNet and SD map prior.
-Method Dist Err. AP
+Method Dist Err AP
 Type Seg Pt ped. div. bou. mean
@@
 Noise 1m 10% 5% 0.18m 61.6 64.3 64.8 63.6
+Noise 1m 10% 5% 0.20m 62.5 66.3 64.5 64.4
 Mask 10% 5% 0.43m 61.8 64.0 64.3 63.4
-Table 3: Ablation experiments of the pre-training strategies on MapTRv2.
-Here, we fixed prior_num=2, search_range=5m and integration=replace.
+Table 5: Ablation of the pre-training strategies.
+Here, we fixed prior_num=multiple, search_range=5m and integration=replace.
 “Seg” and “Pt” represent segment-level and point-level, respectively.
-Main Results
-To evaluate the effectiveness of our method across different model architectures, evaluation metrics, and vector prior maps, we integrated the UVE-encoded prior maps into both HDMapNet and the vectorized model MapTRv2 (see Table 1).
-The prior information provided by SD maps, which primarily consists of a centerline skeleton, resulted in the smallest improvement.
-Specifically, MapTRv2 and HDMapNet showed enhancements of 3.0 mAP and 1.9 mIoU, respectively, when using SD map priors.
-And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior information.
-The incorporation of online local maps, which offer up-to-date information, led to more significant improvements.
-This is reflected in the enhancements of 4.2 mAP and 2.2 mIoU for MapTRv2 and HDMapNet, respectively.
-These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information.
-Figure 4 illustrates how online local maps effectively restore obscured lane dividers, pedestrian crossings, and road boundaries.
-Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively.
-Additionally, we evaluated the results using a stricter threshold of 0.5m.
-Under this criterion, our method with SD map, online local map, and outdated HD map (HD map-EX*) priors achieved improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively.
-These results underscore that our method provides more substantial improvements under stricter evaluation conditions.
-Overall, our method demonstrated consistent performance improvements across all baseline models with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
-Ablation Study
+* indicates data augmentation to expand the data volume.
+4.3 Ablation study
 In this section, we do extensive research on UVE structure, pre-training, number and search range of prior maps, and integration methods.
@@
 However, when UVE was configured with two layers of intra-vector and inter-vector attention and a feature dimension of 64, the model’s performance reached saturation.
+Fixed Prior AP
+Parameters Num ped. div. bou. mean
+- None 59.8 62.4 62.4 61.5
+search_range=5m One 61.1 63.8 63.6 62.8
+pretrain=true Two 61.4 66.2 63.0 63.5
+integration=replace Multiple 62.5 66.3 64.5 64.4
+Table 6: Ablation experiments on the number of online local prior maps on MapTRv2.
+Fixed Search AP
+Parameters Range ped. div. bou. mean
+- None 59.8 62.4 62.4 61.5
+prior_num=2 5 m 62.5 66.3 64.5 64.4
+pretrain=true 10 m 61.4 64.6 64.4 63.5
+integration=replace 15 m 61.9 63.2 63.9 63.0
+Table 7: Ablation experiments on the search range of online local prior maps on MapTRv2.
 The effect of pre-training.
-To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies, as shown in Table 3.
+To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies in Table 5.
 When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters.
-However, the simplicity of the task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement.
+However, the simple task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement.
 We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly.
-The UVE’s average Euclidean distance error in reconstructing the entire map was 0.46 meters.
+The UVE’s average Euclidean distance error was 0.46 meters.
 However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged.
@@
 We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement.
-This suggests that while challenging tasks based on masking can pre-train UVE effectively, their impact may be constrained by the amount of pre-training data.
+Although various proposed tasks can effectively pre-train UVE, they are still limited by the amount of pre-training data.
+Therefore, we employ horizontal and vertical flipping data augmentation to expand the pre-training dataset.
+The result of the data augmentation is shown as *, effectively improving 1.4 mAP.
+Fixed Integrated AP
+Parameters Method ped. div. bou. mean
+- None 59.8 62.4 62.4 61.5
+prior_num=2 Replace 62.5 66.3 64.5 64.4
+pretrain=true Add 63.3 64.1 64.4 63.9
+search_range=5m Concat 63.6 67.3 66.2 65.7
+Table 8: Ablation experiments on the integrated methods of online local prior maps on MapTRv2.
 Impact of the number of prior maps.
-We conducted an ablation study to assess the effect of varying the number of prior maps on model performance (see Table 4).
-The results indicate that using a single prior map led to a 0.9 mAP improvement, demonstrating its effectiveness in enhancing current perception.
-Incorporating two prior maps resulted in a more substantial improvement of 1.5 mAP, reflecting the value of richer prior information.
-However, using multiple prior maps only yielded a 1.3 mAP improvement.
-This could be attributed to the integration method, where replacing too many original queries with prior features may have diminished the performance of the original queries.
+We conducted an ablation study to assess the effect of the number of prior maps on model performance (see Table 6).
+The results indicate that using a single prior map led to a 1.3 mAP improvement, demonstrating its effectiveness in enhancing current perception.
+Incorporating two prior maps resulted in a more substantial improvement of 2.0 mAP, reflecting the value of richer prior information.
+Finally, multiple prior maps effectively improved 2.9 mAP, indicating that more prior maps can provide more comprehensive necessary information.
 Effect of different search ranges for prior maps.
@@
 To evaluate the versatility of our approach, we tested various integration methods in Table 8.
-Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.1 and 1.6, respectively, but these methods may impact the performance of the original query.
+Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.9 and 2.4, respectively, but these methods may impact the performance of the original query.
 In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP.
@@
 Our proposed UVE accounts for only 15.3% of the total pipeline runtime, with the majority of inference time spent on the 2D Backbone (50.2%), highlighting the efficiency and practical value of our approach across various application scenarios.
-Fixed Prior AP
-Parameters Num ped. div. bou. mean
-- None 59.8 62.4 62.4 61.5
-search_range=5m One 61.9 62.1 62.4 62.4
-pretrain=false Two 61.3 63.4 64.4 63.0
-integration=replace Multiple 61.2 63.0 64.2 62.8
-Table 4: Ablation experiments on the number of online local prior maps on MapTRv2.
-Fixed Search AP
-Parameters Range ped. div. bou. mean
-- None 59.8 62.4 62.4 61.5
-prior_num=2 5 m 61.3 63.4 64.4 63.0
-pretrain=false 10 m 60.8 63.7 64.0 62.8
-integration=replace 15 m 61.5 62.7 64.0 62.7
-Table 5: Ablation experiments on the search range of online local prior maps on MapTRv2.
-Fixed Integrated AP
-Parameters Method ped. div. bou. mean
-- None 59.8 62.4 62.4 61.5
-prior_num=2 Replace 61.6 64.3 64.8 63.6
-pretrain=true Add 62.5 63.3 63.6 63.1
-search_range=5m Concat 63.6 67.3 66.2 65.7
-Table 6: Ablation experiments on the integrated methods of online local prior maps on MapTRv2. Note that we proposed other integrated methods after pre-training UVE, so we use pre-trained UVE here.
 - MapTRv2 Ours
 FPS 10.3 9.9
-GPU mem. (MB) 2624 2650
+GPU mem (MB) 2624 2650
 Params (MB) 40.3 43.4
@@
 2D Backbone 47.6 50.2%
-PV to BEV Module 12.9 13.6%
+PV to BEV 12.9 13.6%
 UVE 14.6 15.3%
@@
 Table 9: Detailed comparison of our method with MapTRv2. FPS and detailed runtime of each component are measured on one A6000 GPU with batch size as 1.
-Figure 4: Qualitative results with and w/o online local prior. Prior maps can help restore obscured map elements and make predictions more complete and accurate.
-Conclusion
-In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction for autonomous vehicles.
-Central to our approach is the UVE, which we designed to encode diverse vector data efficiently.
-Through comprehensive experimentation, we demonstrated that UVE, combined with our proposed pre-training strategies, significantly improves the performance of state-of-the-art mapping models.
-Our approach not only addresses the challenges associated with real-time HD map construction in dynamic and complex environments but also offers a scalable solution that continually improves map accuracy over time.
-The iterative use of historical prediction maps as priors leads to progressively refined map outputs, thus ensuring that the most up-to-date road information is available for autonomous navigation.
+Figure 5: Qualitative results with and w/o online local prior.
+5 Conclusion
+In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction.
+Central to PriorDrive is UVE, which can encode diverse vector data efficiently.
+Through comprehensive experiments, we demonstrated that UVE, combined with proposed pre-training strategies, significantly improves performance of state-of-the-art models in various vector tasks.
+PriorDrive not only addresses the challenges associated with online HD map construction in complex environments but also offers a scalable solution that continually improves accuracy over time.
+The iterative use of historical prediction maps as priors leads to progressively refined map outputs, providing latest road information for autonomous driving.

### v4

@@
-Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping
+PriorDrive: Enhancing Online HD Mapping with Unified Vector Priors
 Abstract
 High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges.
-The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather.
-This paper proposes the PriorDrive framework to addresses these limitations by harnessing the power of prior maps, significantly enhancing the robustness and accuracy of online HD map construction.
-Our approach integrates a variety of prior maps, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data.
-To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements.
-At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs hybrid prior embedding and a dual encoding mechanism to process vector data.
-Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance.
+The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather, while their performance in distant regions remains unsatisfying.
+This paper proposes PriorDrive to address these limitations by directly harnessing the power of various vectorized prior maps, significantly enhancing the robustness and accuracy of online HD map construction.
+Our approach integrates a variety of prior maps uniformly, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data.
+To effectively integrate such prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements.
+We further propose a Unified Vector Encoder (UVE), which employs fused prior embedding and a dual encoding mechanism to encode vector data.
+To improve the UVE’s generalizability and performance, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data.
 Through extensive testing on the nuScenes, Argoverse 2 and OpenLane-V2, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities.
-The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
+The integration of prior maps through PriorDrive offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
+Code is available at https://github.com/MIV-XJTU/PriorDrive
 Figure 1: Overview of PriorDrive.
-Our unified vector encoder (UVE) directly encodes diverse vector prior maps and seamlessly integrates them into existing online mapping frameworks.
-This integration enhances the final predictions, making them more complete and accurate than those generated without prior information.
+PriorDrive seamlessly integrates diverse vectorized prior maps into existing online mapping frameworks, making final predictions more complete and accurate than those generated without priors.
+The optimized predictions can be uploaded to cloud servers for other vehicles to download and use as prior maps.
 footnotetext: Work done during the internship at Amap, Alibaba Group.footnotetext: Equal contribution.footnotetext: Corresponding author.
-1 Introduction
-High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries [6, 8].
+This work was supported by the National Natural Science Foundation of China No.
+62572385 and the Fundamental Research Funds for the Central Universities No.
+xxj032023020.
+Introduction
+High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements (HDmaps-survey; Gao2020VectorNetEH; yao2024waterscenes; yao2025exploring; lu2025differentiable).
 Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive.
 These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments.
-In response, there is increasing interest in online HD map construction [27, 5, 38], where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds.
-Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
-In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps.
-Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure 1.
-SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity.
-Existing HD maps (HD map-EX), despite their high accuracy, often fail to reflect current road conditions due to infrequent updates.
-Nevertheless, these maps offer valuable prior information for HD map construction.
-Additionally, historical prediction maps, which are often overlooked, can provide insights into current road conditions by reflecting the latest observations from previous perceptions.
-As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
-Recent studies [35, 13, 29] have attempted to use prior maps to supplement perception data, but they face several limitations.
-For instance, NMP [35] relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments.
-P-MapNet [13] incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively.
-MapEX [29], on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps.
-These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
-In addition to constructing accurate online HD maps, the autonomous driving scenario requires understanding the topology in road scenes (i.e., lane connectivity and correspondence with traffic elements), which is called topology reasoning task.
-Since connected SD maps include crucial information for road topology, we also apply PriorDrive to this task to enhance the performance of topology reasoning.
-Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model.
-However, designing such a model is challenging due to the variable length, sparsity, and discreteness of vector data.
-To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE).
-The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow model to fully learn the prior distribution of vector maps.
-Our experiments on nuScenes, Argoverse 2 and OpenLane-V2 demonstrate that our PriorDrive is plug-and-play and can be seamlessly applied to various online mapping models and vector tasks, resulting in significant performance improvements.
+In response, there is increasing interest in online HD map construction (BEMAPNET; PivotNet; MAPVR; yao2025usvtrack), where maps are generated in real-time using on-board sensors.
+Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions (see red circle in Figure 1) or inclement weather conditions, coupled with unsatisfactory performance in distant regions.
+Using prior maps to alleviate the problem of online mapping is a promising solution.
+Specifically, there are three common types of prior maps: 1) Standard Definition Maps (SD maps), 2) existing offline outdated HD maps, and 3) online historical prediction local maps, as illustrated in the upper left corner of Figure 1.
+SD maps provide crucial long-distance centerline skeletons, but with uncertain accuracy compared to HD maps.
+While existing HD maps (HD map-EX) offer high accuracy, their infrequent updates may lead to outdated information that fails to fully reflect current road conditions.
+Historical predicted local maps can offer insights by incorporating previous observations, but being a single mapping result, they may not guarantee the completeness and accuracy of the prior map.
+Generally, there are two ways to encode vectors: 1) rasterized, 2) vectorized.
+Recent studies (Xiong2023NeuralMP; Jiang2024PMapNetFM; li2024towards1; li2025hyfacialhybrid) have attempted to use rasterized or vectorized prior maps to alleviate the problem of online mapping.
+However, existing prior-based methods (Xiong2023NeuralMP; Jiang2024PMapNetFM; MAPEX; 10.1117/12.3049486; li2025lion) face several limitations.
+For instance, P-MapNet (Jiang2024PMapNetFM), NMP (Xiong2023NeuralMP) and HRMapNet (zhang2024hrmapnet) leverage rasterized SD maps or historical prediction maps.
+The rasterized representation, limited by resolution, is lossy and redundant while lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively.
+Furthermore, P-MapNet and HRMapNet require complex post-processing to convert vectorized maps into rasterized maps, which also do not align with the native vector storage format of maps.
+On the other hand, MapEX (MAPEX) exclusively relies on vectorized outdated HD maps, which struggle to reflect real-time changes in road structures, due to low update frequencies, making them unreliable for current navigation needs.
+Moreover, the encoding scheme of MapEX is overly simplistic, thereby failing to effectively capture instance-level information within vector representations.
+It is worth mentioning that encoding vectors is not a straightforward task.
+There are several main challenges: 1) Different types of vector maps may contain varying vector types (e.g., points, lines, planes), 2) Vector lengths can be fixed or variable (see blue circle in Figure 1), and 3) Alignment and inconsistency issues exist between different vector maps (see purple circle in Figure 1).
+Previous methods were limited to efficiently encoding only a single type of prior map, but the complementary information contained across different prior map categories.
+For instance, NavMap (schmidt2023navmap) omitted road segments with variable lengths when integrating HD maps and SD maps, resulting in the loss of valuable information.
+Furthermore, previous methods generally encode only vector positions and categorical attributes, lacking the capability to capture finer-grained information such as direction, geometric shape, and topology.
+In this paper, we overcome these challenges and introduce PriorDrive, a novel framework designed to directly and unifiedly integrate diverse vectorized prior maps into various mapping models, fully leveraging their respective strengths and complementary information.
+We propose a Unified Vector Encoder (UVE) that encodes various vector data and captures fine-grained vector information.
+It can extract fixed-length instance-level and point-level features from variable-length vectors.
+Additionally, we introduce for the first time a point-level and segment-level pre-training paradigm specifically tailored for vector data, aiming to enhance the model’s understanding and representation of vector.
+Concurrently, we propose a hybrid prior representation (HPQuery) to represent all elements, facilitating the integration of unified vector features into mapping models.
+Experiments on nuScenes, Argoverse 2 and OpenLane-V2 demonstrate that PriorDrive is plug-and-play, seamlessly applicable to diverse online mapping models and vector tasks with significant performance improvements.
 In summary, our contributions are as follows:
-  • We introduce a unified vector encoder that effectively captures diverse vector attributes and enables encoding of variable-length vectors into fixed-length features through Hybrid prior embedding and dual encoding mechanism.
-  • We first proposed a pre-training paradigm for vector data, which can effectively learn the prior distribution of vector data and help reduce noise in historical maps through point-level and segment-level pre-training.
-  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception.
-Our comprehensive evaluation demonstrates that PriorDrive is plug-and-play and significantly enhances the performance of online mapping models.
-2 Related work
-2.1 Online vectorized HD map construction
-Online vectorized HD map construction was initially treated as a segmentation task [15, 18].
-To construct vectorized HD map, HDMapNet [15] requires complex post-processing of pixel-level rasterized maps.
-VectorMapNet [25] introduced a coarse-to-fine, two-stage network leveraging keypoint representations.
-MapTR [19] advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like [2] one-stage network.
-Subsequent studies [41, 39, 36, 9, 12, 40, 3] have introduced various insightful approaches for further improvement.
-However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
-The release of OpenLane-V2 [31] dataset has sparked growing interest in topology reasoning within driving scenes.
+  • We introduce a unified vector encoder that effectively captures diverse vector attributes and enables encoding of variable-length vectors into fixed-length features through fused prior embedding and dual encoding mechanism.
+  • We first proposed a point-level and segment-level pre-training paradigm for vector data, effectively capturing prior distribution and mitigating noise in historical maps.
+  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with various vector prior maps to address the limitations of single-perception.
+Our comprehensive evaluation demonstrates that PriorDrive is plug-and-play and significantly enhances the online mapping models.
+Related Work
+Online Vectorized HD Map Construction
+Online vectorized HD map construction was initially treated as a segmentation task (bevformer; zhang2024self; sun2023large; guo2025depth; cui2025efficient).
+To construct a vectorized HD map, HDMapNet (Li2021HDMapNetAO) requires complex post-processing of pixel-level rasterized maps.
+VectorMapNet (Liu2022VectorMapNetEV) introduced a coarse-to-fine, two-stage network leveraging keypoint representations.
+MapTR (liao2022maptr) advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like (detr; 9681188; zeng2025janusvln; xiao2025worldenv; jin2025reasoning) one-stage network.
+Subsequent studies (HIMAP; sun2025yolov4svm; zeng2024driving; lan2025contextual; lan2025mappo) have introduced various insightful approaches for further improvement.
+However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion, inclement weather conditions or distant regions.
+The release of OpenLane-V2 (wang2023openlanev2) dataset has sparked growing interest in topology reasoning within driving scenes.
 This task focuses on recognizing the topologies among lanes and lanes with traffic elements.
-TopoNet [17] leverages GNN to update lane representation and lane topology.
-TopoMLP [34] utilizes PETR [24] for centerline detection and employs simple MLPs to predict the relationships.
-In this paper, we also apply PriorDrive to topological inference tasks, introducing interconnected SD Map as an additional input to support topology reasoning.
-2.2 Online mapping based on prior maps
+TopoNet (li2023toponet) leverages GNN to update lane representation and lane topology.
+TopoMLP (wu2023topomlp) utilizes PETR (Liu2022PETRPE) for centerline detection and employs simple MLPs to predict the relationships.
+In this paper, we also apply PriorDrive to topological inference tasks, demonstrating that PriorDrive is plug and play and widely applicable to other vector tasks.
+Online Mapping Based on Prior Maps
 Recent research has explored the integration of prior maps to enhance the performance of online mapping models.
-P-MapNet [13] encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps.
-NMP [35] introduces a global neural map prior that self-updates, thereby improving the performance of local map inference.
-MapEX [29] categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm.
-Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information.
-Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
-2.3 Pretrained methods based on mask modeling
+P-MapNet (Jiang2024PMapNetFM) encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps.
+NMP (Xiong2023NeuralMP) introduces a global neural map prior that self-updates, thereby improving the performance of local map inference.
+MapEX (MAPEX) utilizes existing HD maps and refines the query-based map estimation model’s matching algorithm.
+Although these advancements have achieved considerable results, many studies have overlooked the potential of online historical prediction maps as a source of prior information.
+Moreover, previous methods (zhang2024hrmapnet; Ma2024RoadPainterPA; wan2025driving; yuan2025unimapgen; li2025multi) typically could only utilize a single type of prior map.
+Our work can uniformly encode various types of vector prior maps to enhance current perception data, fully leveraging the complementary advantages among different prior maps.
+Pretrained Methods Based on Mask Modeling
 In the fields of NLP and CV, masked pre-training has proven to be an effective strategy for self-supervised representation learning.
-In NLP, models like BERT [4, 23, 14] use masked language modeling to predict randomly masked tokens within a bidirectional text context.
-Similarly, methods such as MAE [11, 16, 21] mask random patches of input images and reconstruct them based on the remaining unmasked patches.
-Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data.
+In NLP, models like BERT (betr; Liu2019RoBERTaAR; Lan2019ALBERTAL; zeng2025FSDrive; li2023ultrare) use masked language modeling to predict randomly masked tokens within a bidirectional text context.
+Similarly, methods such as MAE (MAE; Liu2022MixMAEMA; liang2025persistent; dai2025unbiased; yin2025knowledge) mask random patches of input images and reconstruct them based on the remaining unmasked patches.
+Diverging from these approaches (zhang2025molebridge; ZHANG2025103888; wei2025copeft; wei2026infocom; li2025cogvla), which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data.
 Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized data.
-3 Proposed method: PriorDrive
-Figure 2: Our Integration Method.
-For query-based models, we enable interaction between the prior features and queries at both instance-level and point-level, ensuring a more nuanced integration.
-Hierarchical query comes from MapTRv2 [20].
-For the other models, we integrate prior features directly with the BEV features.
-3.1 Problem formulation
+Proposed Method: PriorDrive
+Problem Formulation
 Formulation of Online Mapping.
@@
 The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features.
-The map decoder usually comprises a Fully Convolutional Network [28] or a query-based module similar to DETR [2], which outputs either rasterized perception results or vectorized map elements.
+The map decoder usually comprises a Fully Convolutional Network (Shelhamer2014FullyCN; hu2024tmff; 202508.0462; huang2024ar) or a query-based module similar to DETR (detr; xie2025seqgrowgraph; tang2025dissecting; tang2025ocrt), which outputs either rasterized perception results or vectorized map elements.
+Figure 2: Structure of UVE. Fused prior embedding enhances vector representation. Intra-instance and inter-instance attention refine local features and capture global context effectively.
 Let the set of images be denoted as $\{I_{1},\ldots,I_{k}\}$, and the BEV encoder be represented as ${E}_{bev}$. The extraction of BEV features ${f}_{bev}$ can then be expressed as:
@@
 $\mathcal{P}={D}_{map}(f_{bev},~opt({Q})),$ (2)
-where $\mathcal{P}$ represents the predicted map elements, $Q$ denotes the learnable query, and $opt$ indicates that this learnable query is optional, which is required in query-based models but not in other models.
+where $\mathcal{P}$ represents the predicted map elements, $Q$ denotes the learnable query, and optopt indicates that this learnable query is optional, which is required in query-based models but not in other models.
+Architecture of UVE
+Inspired by BERT (betr; he2025social; rs17132267; ren2025digital) on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector elements to sentences (see Figure 2).
 Extraction of Prior Map Features.
 Our proposed UVE serves as a unified vector encoder that can directly encode various vector data information, such as the position ($x$, $y$), direction ($v_{x}$, $v_{y}$), and type of points ($c$).
-Using $p=\left[x,y,v_{x},v_{y},c\right]$ to represent the point, a vector $v$ typically consists of multiple ordered sets of points, $v=\left[p_{1},p_{2},\ldots,p_{n}\right]$, where $n$ varies.
-The map prior $M_{prior}$ consists of multiple vectors $M_{prior}=\{v_{1},v_{2},\ldots,v_{m}\}$, with $m$ being variable.
+Using $p=\left[x,y,v_{x},v_{y},c\right]$ to represent the point, a vector $v$ typically consists of an ordered set of a variable number of points, $v=\left[p_{1},p_{2},\ldots,p_{n}\right]$, where $n$ varies.
+The map prior $M$ consists of a set of vectors $M=\{v_{1},v_{2},\ldots,v_{m}\}$, with $m$ being variable.
+When employing different prior maps, all of prior maps were integrated into a cohesive whole, $M_{prior}=\{M_{1},M_{2},\ldots,M_{t}\}$.
 Using $E_{uve}$ to represent the UVE model, the extraction process of the features of the prior map, denoted as $f_{prior}$, can be represented as:
 $f_{prior}=E_{uve}(M_{prior}).$ (3)
-Online mapping based on prior maps.
-We proposed four methods in total to integrate prior map features into online mapping models.
-For the methods without a learnable query $Q$, such as HDMapNet [15], we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{bev}$.
-Then connect it with the BEV feature $f_{bev}$ and aligned by $Conv$, as shown in Figure 2 left.
-$f_{bev}^{*}=Conv(cat[f_{bev},~{}Deconv(reshape(f_{prior}))]),$ (4)
-here, $f_{bev}^{*}$ represents BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
-For methods based on learnable query $Q$, such as the vectorized model MapTR series [20], we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 2 right.
-Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively.
-The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$.
-So we interact prior features and queries at the instance level and click, respectively.
-The HPQuery of addition, replacement, and concatenation operations can be expressed as:
-$q_{ij}^{*}=q_{i}^{ins}.add(f_{i}^{ins})+q_{j}^{pt}.add(f_{j}^{pt}),$ (5)
-$q_{ij}^{*}=q_{i}^{ins}.replace(f_{i}^{ins})+q_{j}^{pt}.replace(f_{j}^{pt}),$ (6)
-$q_{ij}^{*}=concat[q_{i}^{ins},~{}f_{i}^{ins}]+concat[q_{j}^{pt},~{}f_{j}^{pt}].$ (7)
-We will detail the effectiveness of these feature fusion methods in the experimental section.
-Figure 3: Structure of UVE.
-Hybrid prior embedding enhances vector representation.
-Intra-instance and inter-instance attention mechanisms are employed to refine local features and capture global context effectively.
-3.2 Architecture of UVE
-Inspired by BERT [4] on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 3).
-And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
-Hybrid prior embedding.
-As shown in Figure 3, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE).
-We use the method from [30] for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$.
+Fused Prior Embedding.
+As shown in Figure 2, for the given prior maps $M_{prior}$, we first construct the input for UVE, termed as Fused Prior Embedding (FPE).
+We use the method from (TancikSMFRSRBN20; zeng2025lens; lin2025plan; yang2025unleashing) for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$.
 We concatenate these two embeddings as the point-level embedding.
-To obtain the feature representation of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding.
+To obtain the fixed-length feature of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding.
 Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances.
 To ensure the order of points, we also introduce 2D learnable position embedding.
-Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
+Finally, these embeddings are aggregated to form FPE.
+Figure 3: Pre-trained Pipeline. The noise & mask generator creates segment-level and point-level noise or mask, and then reconstructs the entire map via UVE and MLP.
 Dual Encoding Mechanism.
-Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector.
-As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances.
-Figure 4: Pre-trained Pipeline. The noise & mask generator creates segment-level and point-level noise or mask, and then reconstructs the entire map via UVE and MLP.
-3.3 Pre-training UVE: Position modeling
-Due to the limited inference ability of online mapping model, there are errors in historical prediction maps.
+Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector.
+Therefore, UVE encoder utilizes $M$-layer intra-vector attention and $N$-layer inter-vector attention, using an attention masking mechanism to facilitate feature interactions within and between different vector instances.
+Each mask in the intra-vector encoder has a value of 1 at its own position and 0 elsewhere, while the inter-vector encoder employs a mask matrix filled entirely with 1s.
+When using different prior maps and multiple online local maps, there are alignment problems between different prior maps and potential inconsistencies between online local maps generated across different trips or different timestamps in one trip.
+Our UVE mitigated this challenge by employing fused prior embedding and dual encoding mechanism to enhance the model’s ability to learn instances, supporting adaptive focus on the superior map elements in different prior maps.
+The experiments presented in Table 4 substantiate this capability.
+Pre-training UVE: Position Modeling
+Due to the limited inference ability of online mapping models, there are errors in historical prediction maps.
 So we used position modeling to pre-train UVE to improve its encoding and noise reduction capabilities.
@@
 For segment-level noise, random noise is added to all vector points within a sub-vector segment, after randomly selecting $10\%$ of the map elements.
-Using $M_{org}$ to represent the original map elements (with the same data structure as $M_{prior}$), the process of adding Gaussian noise can be represented by the following formula:
+Using $M_{org}$ to represent the original map elements (with the same data structure as $M$), the process of adding Gaussian noise can be represented by the following formula:
 $M_{org}^{*}=M_{org}\emph{[random\_index]}+\epsilon,$ (4)
 where $\epsilon\sim\mathcal{N}(0,1)$ is a random noise, which is only added to the horizontal and vertical coordinates of each point.
-Similar to adding noise to the points, the selection of points for adding the mask is also divided into point-level and segment-level.
-After selecting the points to add the mask, the coordinates of these points will be masked as $-1$.
-[…]
+Similar to adding noise, the selection of points for adding the mask is also divided into point-level and segment-level.
+After selecting the points to add the mask, the coordinates of these points will be masked as 1-1.
+$M_{org}\emph{[random\_index]}=Mask.$ (5)
 Loss Function.
-We feed the vector map to UVE encoding after passing through the noise mask generator, then decode the coordinates of all vector points using MLP, and calculate mean euclide distance from real coordinates as supervision:
+We feed vector map to UVE encoding after passing through the noise mask generator, then decode the coordinates of all points using MLP, and calculate mean euclide distance from GT coordinates as supervision:
 $\mathcal{L}=RMSE(P,~mlp(E_{uve}(M_{org}^{*}))),$ (6)
-where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in $M_{org}$.
-4 Experiments
-4.1 Experimental settings
-Method Modality Backbone Epoch Metric ped. div. bou. mean $\mathbf{\tau_{0.5m}}$
-HDMapNet [15] C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
-# SD map Prior (Ours) C Effi-B0 30 IoU 22.2 42.0 41.5 35.2 -
-# Online Local Map Prior (Ours) C Effi-B0 30 IoU 23.1 41.4 42.0 35.5 -
-# HD map-EX* Prior (Ours) C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
-MapTR [19] C R50 24 AP 41.2 49.5 51.1 47.3 -
-# P-MapNet [13] C R50 24 AP 43.7 50.9 53.5 49.4 -
-StreamMapNet [37] C R50 24 AP 60.4 61.9 58.9 60.4 -
-# SQD-MapNet [32] C R50 24 AP 63.0 65.5 63.3 63.9 -
-MGMap [22] C R50 24 AP 61.8 65.0 67.5 64.8 -
-MapTRv2 [20] C R50 24 AP 59.8 62.4 62.4 61.5 38.6
-# SD map Prior (Ours) C R50 24 AP 61.7 65.5 66.4 64.5 42.7
-# Online Local Map Prior (Ours) C R50 24 AP 63.6 67.3 66.2 65.7 44.7
-# HD map-EX* Prior (Ours) C R50 24 AP 61.5 65.6 68.3 65.1 44.3
-Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range.
-“C” denotes camera input.
-The SD map is sourced from OpenStreetMap [10].
-The online local map refers to the historical prediction results.
-As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX [29] to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
+where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the points in $M_{org}$, $RMSE$ on behalf of the root mean square error.
+Figure 4: Our Integration Method.
+For query-based models, we enable interaction between the prior features and queries at both instance-level and point-level, ensuring a more nuanced integration.
+For the other models, we integrate prior features directly with the BEV features.
+Integrated Approaches to BEV and HPQuery
+We respectively proposed different integration methods for the query-based/non-query-based approaches to integrate the prior map features into the online map construction model.
+For methods without a learnable query $Q$, such as HDMapNet (Li2021HDMapNetAO), we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{bev}$.
+Then connect it with the BEV feature $f_{bev}$ and aligned by $Conv$, as shown in Figure 4 left.
+$f_{bev}^{*}=Conv(cat[f_{bev},~Deconv(reshape(f_{prior}))]),$ (7)
+here, $f_{bev}^{*}$ represents BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
+For methods based on learnable query $Q$, such as the vectorized model MapTR series (maptrv2), we propose HPQuery of three operations: addition, replacement, and concatenation, as shown in Figure 4 right.
+Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively.
+The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$.
+So we interact prior features and queries at the instance-level and point-level, respectively.
+The HPQuery of addition, replacement, and concatenation operations can be expressed as:
+$q_{ij}^{*}=q_{i}^{ins}.add(f_{i}^{ins})+q_{j}^{pt}.add(f_{j}^{pt}),$ (8)
+$q_{ij}^{*}=q_{i}^{ins}.replace(f_{i}^{ins})+q_{j}^{pt}.replace(f_{j}^{pt}),$ (9)
+$q_{ij}^{*}=concat[q_{i}^{ins},~f_{i}^{ins}]+concat[q_{j}^{pt},~f_{j}^{pt}].$ (10)
+We will detail the effectiveness of these feature fusion methods in supplementary material.
+Experiments
+Experimental Settings
 Benchmark.
-We mainly validate our method on the popular nuScenes [1] and Argoverse 2 [33] following the standard setting of previous methods [25, 19].
-NuScenes contains a total of 1,000 scenes, each with an approximate duration of 20 seconds.
-The data was collected utilizing a 32-beam LiDAR operating and six cameras offering a 360-degree field of view.
-Argoverse 2 includes 1000 logs, each offering 15 seconds of 20Hz RGB images captured from 7 cameras, along with a 3D vector map.
-Furthermore, to demonstrate that PriorDrive is plug-and-play and widely applicable to other vector tasks, we conducted a topology inference task on OpenLane-V2 [31] dataset.
-Subset A, developed by Argogorse 2 [33], includes seven views as input, along with an additional SD map.
+We mainly validate our method on the popular nuScenes (Caesar2019nuScenesAM) and Argoverse 2 (Argoverse2) following the previous methods (zhang2024hrmapnet; maptrv2; zhou2025hademif; zhou2025valuing).
+NuScenes contains a total of 1,000 scenes, which was collected utilizing a 32-beam LiDAR operating and six cameras offering a 360-degree field of view.
+Argoverse 2 includes 1000 logs, which was captured from 7 cameras, along with a 3D vector map.
+To demonstrate PriorDrive is plug-and-play and widely applicable to other vector tasks, we conducted topology inference on OpenLane-V2 (wang2023openlanev2) Subset A.
 Evaluation Protocol.
-For the vectorized results, we follow the standard metric used in previous works [15, 25, 19].
-The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis.
-And we adopt average precision (AP) evaluated the quality of map construction.
-We compute the $\mathrm{AP}_{\tau}$ under several Chamfer distance $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
+For vectorized results, we follow the standard metric used in previous works (Liu2022VectorMapNetEV; maptrv2; li2025revolutionizing; li2025knowledge).
+The perception range along the direction of vehicle travel is [15.0m,15.0m][-15.0m,15.0m] for $X$-axis and [30.0m,30.0m][-30.0m,30.0m] for $Y$-axis.
+We compute the average precision (AP) $\mathrm{AP}_{\tau}$ under several Chamfer distance $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the final AP metric:
 $\mathrm{AP}=\frac{1}{|\mathrm{T}|}\sum_{\tau\in\mathrm{T}}\mathrm{AP}_{\tau}.$ (11)
-Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works [15, 35, 13] and employ IoU as the metrics for segmentation results.
+Method Prior Map Backbone Epoch Metric Ped. Div. Bou. Mean
+HDMapNet [ICRA21] (Li2021HDMapNetAO) $\times$ Effi-B0 30 IoU 18.7 40.6 39.5 32.9
+P-MapNet [RAL24] (Jiang2024PMapNetFM) $\checkmark$ Effi-B0 30 IoU 22.6 44.1 43.8 36.8
+NMP [CVPR23] (Xiong2023NeuralMP) $\checkmark$ Effi-B0 30 IoU 21.0 44.2 46.1 37.1
+PriorDrive (Ours) $\checkmark$ Effi-B0 30 IoU 25.1 46.7 48.4 40.1
+MapTR [ICLR23] (liao2022maptr) $\times$ R50 24 AP 41.2 49.5 51.1 47.3
+P-MapNet [RAL24] (Jiang2024PMapNetFM) $\checkmark$ R50 24 AP 43.7 50.9 53.5 49.4
+StreamMapNet [WACV24] (Yuan_2024_streammapnet) Temporal R50 24 AP 60.4 61.9 58.9 60.4
+PriorDrive (Ours) $\checkmark$ R50 24 AP 60.9 65.3 63.8 63.3
+MapTRv2 [IJCV24] (maptrv2) $\times$ R50 24 AP 59.8 62.4 62.4 61.5
+PrevPredMap [WACV25] (peng2024prevpredmap) $\checkmark$ R50 24 AP 64.5 66.9 67.6 66.3
+HRMapNet [ECCV24] (zhang2024hrmapnet) $\checkmark$ R50 24 AP 65.8 67.4 68.5 67.2
+HRMapNet [ECCV24] (zhang2024hrmapnet) $\checkmark$ R50 110 AP 72.0 72.9 75.8 73.6
+InteractionMap [CVPR25] (wu2025interactionmap) Temporal R50 24 AP 69.7 72.7 73.0 71.8
+InteractionMap [CVPR25] (wu2025interactionmap) Temporal R50 110 AP 75.3 75.6 77.0 76.0
+PriorDrive (Ours) $\checkmark$ R50 24 AP 71.8 81.4 74.1 75.8
+PriorDrive (Ours) $\checkmark$ R50 110 AP 78.0 82.0 81.6 80.5
+Table 1: Performance of various online mapping models with unified prior maps on nuScenes.
+All methods use camera input.
+In each block, the first row represents the baseline model, and all other methods are modifications upon the baseline.
+“Temporal” means using temporal information.
+The results of StreamMapNet and HRMapNet are from HRMapNet, while the other results are from the respective papers.
+Method Epoch $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
+MapTRv2 [IJCV24] 6 60.7 68.9 64.5 64.7
+HRMapNet [ECCV24] 30 65.1 71.4 68.6 68.3
+PrevPredMap [WACV25] 6 64.1 71.4 67.4 67.6
+InteractionMap [CVPR25] 6 66.6 75.6 72.7 71.6
+PriorDrive (Ours) 6 68.3 75.8 74.2 72.8
+Table 2: Performance comparison of 3D map elements on Argoverse 2.
+Method Prior Map $\textbf{DET}_{l}$ $\textbf{DET}_{t}$ v1.0.0 v2.1.0
+$\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS $\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS
+TopoNet + OSMR [IROS24] (osmr) $\checkmark$ 30.6 44.6 7.7 22.9 37.7 - - -
+RoadPainter [ECCV24] (Ma2024RoadPainterPA) $\checkmark$ 36.9 47.1 12.7 25.8 42.6 - - -
+SMERF [ICRA24] (SMERF) $\checkmark$ 33.4 48.6 7.5 23.4 39.4 15.4 25.4 42.9
+TopoLogic [NeurIPS24] (fu2024topologic) $\times$ 29.9 47.2 18.6 21.5 41.6 23.9 25.4 44.1
+PriorDrive (Ours) $\checkmark$ 31.5 48.4 24.6 30.7 46.2 31.4 33.7 48.5
+Table 3: Performance comparison with SOTA non-prior/prior-based methods on OpenLane-V2 subset_A set. ’-’ denotes the absence of relevant data.
+For the rasterized results, we follow the previous works (Xiong2023NeuralMP; 202506.2087; chen2025visrl; chen2025sifthinker) and employ IoU as the metric for segmentation results.
 For OpenLane-V2’s lane centerline perception task, $DET_{l}$ averages the Frechet distance to quantify similarity.
@@
 Implementation Details.
-For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes.
-All experiments were trained with 6 NVIDIA RTX A6000.
-For pre-training UVE, we trained the 24-epoch on nuScenes vecror data.
-For historical prediction maps, we run each base model to get their respective online local maps.
+For fair comparisons, experiments were conducted on all baseline models using their original hyperparameters.
+All experiments were trained with 8 NVIDIA RTX A6000.
+For pre-training UVE, we trained the 24-epoch on nuScenes vector data.
+This one-time offline pre-training ( 12h) yields significant improvements, and its cost is amortized through plug-and-play reuse across various models.
+The sources of the three types of prior maps are as follows: The SD map is from OpenStreetMap (Haklay2008OpenStreetMapUS); as for the existing HD prior maps, we follow MapEX (MAPEX).
+For each map element localization, we add noise from a Gaussian distribution with a standard deviation of 1 meter.
+This has the effect of applying a uniform translation to each map element (dividers, boundaries, crosswalks); as for the predicted online local maps, in actual scenarios, the online local maps are generated by vehicles passing through the same area, and their data is stored on cloud servers for other vehicles to retrieve in the future.
+For datasets collected by a single vehicle such as nuScenes, Argoverse 2, and OpenLane-V2, certain road sections are captured through multiple drives.
+Therefore, we use each baseline model to infer the entire dataset, thereby generating online local prior maps at different trips or different times in the same trip, simulating this actual process.
+Due to the different occlusion conditions of different trips, the online local prior map contains more abundant supplementary information for the current driving scene.
+So we first search the historical prediction of different trips, and then search the historical prediction of the past time in the same trip as the local priors.
 Method $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
-MGMap [22] 52.8 67.5 68.1 62.8
-MapQR [26] 60.1 71.2 66.2 65.9
-MapTRv2 [20] 60.7 68.9 64.5 64.7
-# SD Prior (Ours) 62.0 70.2 66.3 66.2
-# Online Local Prior (Ours) 63.9 71.7 68.6 68.1
-# HD map-EX* (Ours) 61.5 69.8 69.4 66.9
-Table 2: Comparison on Argoverse 2 where the 3D map elements are predicted directly.
-4.2 Main results
+PriorDrive 71.8 81.4 74.1 75.8
+     w/o UVE 67.3 74.0 67.8 69.7
+     w/o Pretrain 69.4 76.5 69.6 71.5
+     w/o SD Prior 69.5 77.1 72.2 72.9
+     w/o Online Local 68.4 73.9 69.5 70.6
+     w/o HD Prior 67.3 72.0 68.6 69.3
+Table 4: The ablation of each component of PriorDrive.
+Main Results
 Results on nuScenes.
-To evaluate the effectiveness of our method across different model, evaluation metrics, and vector prior maps, we choose HDMapNet and MapTRv2 as baselines (see Table 1).
-SD maps, which primarily consists of a centerline skeleton, showed enhancements of 3.0 mAP for MapTRv2 and 2.3 mIoU for HDMapNet, respectively.
-And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior maps.
-Online local maps, which offer up-to-date information, led to more significant improvements, with 4.2 mAP and 2.6 mIoU for MapTRv2 and HDMapNet, respectively.
-These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information.
-Figure 5 illustrates how online local maps effectively restore obscured map elements.
-More visualization are shown in the supplementary material.
-Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively.
-Additionally, we evaluated the results using a stricter threshold of 0.5m, achieving improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively, underscoring that our method provides more substantial improvements under stricter evaluation conditions.
-Overall, our method demonstrated consistent performance improvements across all baselines with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
-Method $\textbf{DET}_{l}$ $\textbf{DET}_{t}$ $\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS
-MapTR [19] 17.7 43.5 5.9 15.1 31.0
-TopoNet [17] 28.6 48.6 10.9 23.8 39.8
-TopoLogic [7] 29.9 47.2 23.9 25.4 44.1
-# SD Prior (Ours) 29.9 47.1 29.6 32.9 47.2
-# Online Local Prior (Ours) 31.0 47.8 30.9 32.5 47.9
-Table 3: Performance comparison with SOTA methods on OpenLaneV2 subset_A set, where SD Prior comes from the SD map of the openLanev2 itself.
-Since the map elements here have only centerline and no boundaries, there is no HD map-EX* prior.
+We evaluate our method across different model, evaluation metrics in Table 1.
+Our PriorDrive achieved a 7.2 mIoU improvement over baseline HDMapNet, which demonstrates that our vector prior remains effective even for rasterized models and metrics.
+In single-variable comparisons with prior/temporal-based methods employing MapTR as baseline, PriorDrive has demonstrated the most substantial improvement of 16 mAP.
+We further benchmarked our approach against the mainstream MapTRv2 baseline, which demonstrated an average performance gain greater than those achieved on HDMapNet, demonstrating that our HPQuery achieves more comprehensive and fine-grained exploitation of prior information.
+Furthermore, PriorDrive outperforms existing prior/temporal-based SOTA methods that adopt the MapTRv2 baseline, demonstrating the effectiveness of PriorDrive in fully utilizing complementary information from different prior maps.
+Figure 5 shows that using the prior maps can effectively recover the occluded elements, as well as more accurate mapping results.
+Overall, our method demonstrated consistent performance improvements across all baselines, highlighting its generalizability and potential applicability to other frameworks.
+Figure 5: Comparison of qualitative results between those with and without unified prior maps.
 Results on Argoverse 2.
-In Table 2, we present the results of Argoverse 2, which provides 3D vectorized map elements with additional height information as ground truth.
-These testings are all trained 6 epochs with ResNet50 as the backbone.
-Our PriorDrive improves performance by 1.5, 3.4, and 2.2, respectively, using different prior maps, demonstrating the superior generalization ability of PriorDrive in terms of 3D vectorized map construction.
+In Table 2, we present the results of 3 dim maps on Argoverse 2.
+PriorDrive improves performance by 8.1 mAP and outperforms the previous prior/temporal-based SOTA methods, demonstrating excellent generalization performance.
 Results on OpenLane-V2.
-We compared PriorDrive with SOTA methods when using ResNet-50 for training 24 epochs.
-Table 3 shows the results on subset A of OpenLane-V2.
-Compared to TopoLogic, our method achieves better topological reasoning accuracy using connected SD priors (29.6 v.s.
-23.9 on $TOP_{ll}$, 32.9 v.s.
-25.4 on $TOP_{lt}$), and the overall score OLS increased by 3.1, indicating that connected SD is more beneficial for lane connectivity and correspondence with traffic elements.
-Employing online local priors resulted in decent detection accuracy for both detection and topological reasoning, with OLS rising to 47.9.
+We compared PriorDrive on OpenLane-V2 when using ResNet-50 for training 24 epochs in Table 3.
+Compared to TopoLogic baseline, our method achieves better topological reasoning accuracy.
+And our overall score OLS reached 46.2 (v1.0.0) and 48.5 (v2.1.0), outperforming both Topologic and previous prior-based approaches.
 Overall, PriorDrive is a plug-and-play solution that can be widely applied to other vector tasks.
-M N Dim IoU
+Ablation Study
+we conducted ablation experiments on nuScenes based on MapTRv2 unless otherwise specified. More experiments are included in supplement.
+Ablation Study of PriorDrive.
+The ablation study in Table 4 investigates the contribution of each component.
+Specifically, replacing the UVE with separate MLPs for each prior degrades performance by 6.1 mAP, demonstrating UVE’s effectiveness in mitigating the inconsistency and weak alignment among different priors.
+indicating it can also alleviate these issues and enhance generalization.
+The performance degradation observed when ablating any single prior highlights that they contain complementary information, which PriorDrive is the first to leverage simultaneously.
+Method $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
+StreamMapNet 29.6 30.1 41.9 33.9
+HRMapNet 36.9 30.3 44.0 37.1
+MapTracker 45.9 30.0 45.1 40.3
+PriorDrive (Ours) 44.7 35.5 47.3 42.5
+Table 5: Comparisons on the new nuScenes split data sets proposed in StreamMapNet.
+Comparison on New Split.
+The original split of nuScenes exhibits geographical overlap between training and validation sets.
+Therefore, we adopt new split proposed by StreamMapNet in Table 5, which training and validation data are separated in locations.
+PriorDrive achieves 42.5 mAP, which outperforms the previous SOTA methods.
+Method Range $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
+[…]
+StreamMapNet 37.2 42.3 30.2 36.6
+HRMapNet 43.1 47.7 34.9 41.9
+PriorDrive (Ours) 45.0 49.5 37.3 43.9
+[…]
+StreamMapNet 22.4 31.3 16.0 23.3
+HRMapNet 29.4 34.2 19.7 27.8
+PriorDrive (Ours) 28.8 36.3 21.6 28.9
+Table 6: Performance comparison of large perception ranges on nuScenes.
+Comparison of Large Perception Ranges.
+Existing online mapping methods are constrained in prediction range due to the low resolution of distant areas in the captured images which leads to a significant decline in performance.
+However, PriorDrive leverages various prior maps to enhance accuracy in long-range predictions in Table 6, demonstrating an outstanding foresight capability.
+Conclusion
+In this paper, we introduced PriorDrive, a novel framework that effectively and uniformly leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction.
+Through comprehensive experiments, we demonstrated that UVE, combined with proposed pre-training strategies, significantly improves performance of state-of-the-art models in various vector tasks.
+PriorDrive not only addresses the challenges associated with online HD map construction in complex environments but also offers a scalable solution that continually improves accuracy over time.
+The iterative use of historical prediction maps as priors leads to progressively refined map outputs, providing latest road information for autonomous driving.
+Supplementary Material
+Appendix A More Experiments
+Here, we present a detailed examination of additional ablation studies on MapTRv2 unless otherwise specified.
+M N Dim AP
 Intra Inter Ped. Div. Bou. Mean
-4 4 64 20.7 41.3 41.1 34.4
-6 2 64 20.5 41.1 40.9 34.2
-2 6 64 20.7 41.0 41.3 34.3
-2 2 128 19.5 41.4 41.2 34.0
-2 2 64 20.8 41.7 41.8 34.8
-1 1 32 19.8 41.1 40.9 33.9
-Table 4: Ablation experiments on UVE structure within HDMapNet and SD map prior.
-Method Dist Err AP
+w/o UVE 67.3 74.0 67.8 69.7
+4 4 64 71.3 80.0 74.2 75.2
+2 2 128 71.0 80.2 73.5 74.9
+2 2 64 71.8 81.4 74.1 75.8
+1 2 64 70.2 79.1 73.2 74.2
+2 1 64 70.4 78.5 71.9 73.6
+1 1 32 68.1 76.3 69.6 71.3
+Table 7: Ablation experiments of UVE structure.
+Structure of UVE.
+Table 7 presents the results of ablation experiments on the UVE structure.
+As the UVE model parameters are progressively augmented, the mAP increases to 75.8, while the model maintains a lightweight architecture.
+However, performance saturates with further parameter increases, constrained by the limited volume and diversity of the vector data.
+Method AP
 Type Seg Pt Ped. Div. Bou. Mean
-Without Pretrain - 61.3 63.4 64.4 63.0
-Original - - 0.09m 61.2 64.1 64.5 63.3
-Noise 3m 20% - 0.46m 61.4 63.0 63.8 62.7
-Noise 1m 10% - 0.15m 61.6 63.8 64.2 63.2
-Noise 1m 10% 5% 0.18m 61.6 64.3 64.8 63.6
-Noise 1m 10% 5% 0.20m 62.5 66.3 64.5 64.4
-Mask 10% 5% 0.43m 61.8 64.0 64.3 63.4
-Table 5: Ablation of the pre-training strategies.
-Here, we fixed prior_num=multiple, search_range=5m and integration=replace.
-“Seg” and “Pt” represent segment-level and point-level, respectively.
-* indicates data augmentation to expand the data volume.
-4.3 Ablation study
-In this section, we do extensive research on UVE structure, pre-training, number and search range of prior maps, and integration methods.
-we conduct experiments using online local map priors on MapTRv2 unless otherwise specified.
-Structure of UVE.
-Table 4 presents the results of ablation experiments on the UVE structure using the SD prior map within HDMapNet.
-We observed that increasing the number of layers or expanding the feature dimension of UVE yielded only marginal improvements, averaging around 1.3 mIoU.
-This limited gain may be attributed to the model’s excess capacity, leading to overfitting.
-Conversely, a minimal configuration of UVE, with only one layer each of intra-vector and inter-vector attention and a dimension of 32, resulted in a modest improvement of just 1 mIoU.
-However, when UVE was configured with two layers of intra-vector and inter-vector attention and a feature dimension of 64, the model’s performance reached saturation.
+w/o Pretrain 69.4 76.5 69.6 71.5
+Original - - 69.9 77.1 70.2 72.4
+Noise 3m 20% - 69.8 78.0 71.7 73.2
+Noise 1m 10% - 71.0 79.2 72.7 74.3
+Noise 1m 10% 5% 72.0 80.1 73.5 75.2
+Mask 10% 5% 71.4 79.8 73.5 74.9
+Mask & Noise 1m 10% 5% 71.8 81.4 74.1 75.8
+Table 8: Ablation of the pre-training strategies. “Seg” and “Pt” represent segment-level and point-level, respectively.
+The Effect of Pre-training.
+To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies in Table 8.
+When we fed the original vector map directly to UVE without any noise, the simple task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.9 mAP improvement.
+We randomly add segment-level or point-level Gaussian noise to map elements.
+Among them, 10% segment-level noise and 5% point-level noise increase 3.7 mAP the most.
+This demonstrates that strategy incorporating both segment-level and point-level noise are more effective at learning the prior information.
+We also experimented with masking instead of noise, which resulted in a 3.4 mAP improvement.
+Finally, we conducted pre-training by simultaneously applying gaussian noise and mask, achieving the best result of 75.8 mAP.
 Fixed Prior AP
 Parameters Num Ped. Div. Bou. Mean
-- None 59.8 62.4 62.4 61.5
-search_range=5m One 61.1 63.8 63.6 62.8
-pretrain=true Two 61.4 66.2 63.0 63.5
-integration=replace Multiple 62.5 66.3 64.5 64.4
-Table 6: Ablation experiments on the number of online local prior maps on MapTRv2.
+w/o online local None 68.4 73.9 69.5 70.6
+search_range=5m One 70.6 78.6 71.7 73.6
+pretrain=true Two 70.6 79.2 72.4 74.1
+integration=replace Multiple 71.8 81.4 74.1 75.8
+Table 9: Ablation study on the number of online local prior maps.
+Impact of the Number of Prior Maps.
+In history, different vehicles passed through the same area.
+Due to the varying degrees of obstruction, these vehicles would predict different online local maps, thus being able to be used as prior maps.
+Also, the online local maps predicted by the same vehicle a few seconds ago also have reference value, so they can also be used as prior maps.
+Therefore, we conducted an ablation study to assess the effect of the number of online local prior maps (see Table 9).
+The results indicate that using a single online local prior map led to a 3.0 mAP improvement, demonstrating its effectiveness in enhancing current perception.
+Incorporating two prior maps resulted in a more substantial improvement of 3.5 mAP, reflecting the value of richer prior information.
+Due to the varying number of available online local prior maps in different regions, and because of the hyperparameter settings of UVE that can encode vector points, the number of multiple online local prior maps is generally 3 to 5.
+Multiple prior maps effectively improved 5.2 mAP, indicating that our UVE can adaptively alleviate potential inconsistencies across different online local maps and more prior maps can provide more comprehensive necessary information.
 Fixed Search AP
 Parameters Range Ped. Div. Bou. Mean
-- None 59.8 62.4 62.4 61.5
-prior_num=2 5 m 62.5 66.3 64.5 64.4
-pretrain=true 10 m 61.4 64.6 64.4 63.5
-integration=replace 15 m 61.9 63.2 63.9 63.0
+w/o online local None 68.4 73.9 69.5 70.6
+prior_num=2 5 m 71.8 81.4 74.1 75.8
+pretrain=true 10 m 70.2 79.4 72.7 74.1
+integration=replace 15 m 69.0 76.7 70.8 72.2
 Table 10: Ablation experiments on the search range of online local prior maps on MapTRv2.
-The effect of pre-training.
-To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies in Table 5.
-When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters.
-However, the simple task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement.
-We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly.
-The UVE’s average Euclidean distance error was 0.46 meters.
-However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged.
-Next, we applied 1-meter segment-level noise to 10% of the map elements randomly, resulting in a reduced error of 0.15 meters for UVE and a 0.2 mAP improvement compared to the model without pre-training.
-Building on this, we introduced 1-meter point-level noise to 5% of the vector points.
-Although this increased UVE’s error to 0.18 meters, it yielded a more substantial mAP improvement of 0.6.
-This demonstrates that pre-training strategies incorporating both segment-level and point-level noise are more effective at learning the prior information of vector maps.
-We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement.
-Although various proposed tasks can effectively pre-train UVE, they are still limited by the amount of pre-training data.
-Therefore, we employ horizontal and vertical flipping data augmentation to expand the pre-training dataset.
-The result of the data augmentation is shown as *, effectively improving 1.4 mAP.
+Effect of different search ranges for prior maps.
+We search for available online local prior maps based on the distance between the geographical location when the online local prior map was generated and the current geographical location of the vehicle.
+So we performed ablation experiments to explore the impact of search ranges for prior maps in Table 10.
+The results show that as the search range is narrowed, the effect gradually improves.
+Although a smaller search range may reduce the number of online local prior maps, it will be more accurate in terms of geographical overlap.
+In such a situation of slight inconsistency and misalignment, our PriorDrive can effectively alleviate the problem.
 Fixed Integrated AP
 Parameters Method Ped. Div. Bou. Mean
-- None 59.8 62.4 62.4 61.5
-prior_num=2 Replace 62.5 66.3 64.5 64.4
-pretrain=true Add 63.3 64.1 64.4 63.9
-search_range=5m Concat 63.6 67.3 66.2 65.7
+w/o online local None 68.4 73.9 69.5 70.6
+prior_num=2 Replace 71.5 79.6 73.1 74.7
+pretrain=true Add 71.8 80.2 73.6 75.2
+search_range=5m Concat 71.8 81.4 74.1 75.8
 Table 11: Ablation experiments on the integrated methods of online local prior maps on MapTRv2.
-Impact of the number of prior maps.
-We conducted an ablation study to assess the effect of the number of prior maps on model performance (see Table 6).
-The results indicate that using a single prior map led to a 1.3 mAP improvement, demonstrating its effectiveness in enhancing current perception.
-Incorporating two prior maps resulted in a more substantial improvement of 2.0 mAP, reflecting the value of richer prior information.
-Finally, multiple prior maps effectively improved 2.9 mAP, indicating that more prior maps can provide more comprehensive necessary information.
-Effect of different search ranges for prior maps.
-We performed ablation experiments to explore the impact of search ranges for prior maps in Table 7.
-The results indicate that increasing the search range leads to a slight decrease in improvement.
-This suggests that while a larger search range allows the prior map to cover a wider area, it also introduces greater positional error.
-On the other hand, a more precise, narrower search range may not always provide multiple prior maps for most locations, limiting the benefits.
 Different methods for integrating prior maps.
 To evaluate the versatility of our approach, we tested various integration methods in Table 11.
-Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.9 and 2.4, respectively, but these methods may impact the performance of the original query.
-In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP.
+Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 4.1 and 4.6, respectively, but these methods may impact the performance of the original query.
+In contrast, the concatenation method achieved a more substantial improvement of 5.2 mAP.
 This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
-Inference speed, memory, model size, and detailed runtime.
-Table 9 compares our method with MapTRv2 in the above aspects.
-The results show that our method maintains similar FPS, GPU memory usage, and parameter count compared to MapTRv2, while achieving a 4.2 mAP improvement.
-Additionally, we provide the detailed runtime of each component.
-Our proposed UVE accounts for only 15.3% of the total pipeline runtime, with the majority of inference time spent on the 2D Backbone (50.2%), highlighting the efficiency and practical value of our approach across various application scenarios.
 MapTRv2 Ours
@@
 Params (MB) 40.3 43.4
-mAP 61.5 65.7
+mAP 61.5 75.8
 Component Runtime (ms) Proportion
@@
 Map Decoder 19.8 20.9%
-Table 9: Detailed comparison of our method with MapTRv2. FPS and detailed runtime of each component are measured on one A6000 GPU with batch size as 1.
-Figure 5: Qualitative results with and w/o online local prior.
-5 Conclusion
-In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction.
-Central to PriorDrive is UVE, which can encode diverse vector data efficiently.
-Through comprehensive experiments, we demonstrated that UVE, combined with proposed pre-training strategies, significantly improves performance of state-of-the-art models in various vector tasks.
-PriorDrive not only addresses the challenges associated with online HD map construction in complex environments but also offers a scalable solution that continually improves accuracy over time.
-The iterative use of historical prediction maps as priors leads to progressively refined map outputs, providing latest road information for autonomous driving.
+Table 12: Detailed comparison of our method with MapTRv2.
+Inference Speed, Memory, Model Size, and Runtime.
+The results in Table 12 measured on a single RTX A6000 show that our method maintains similar FPS, GPU memory usage, and parameter count compared to MapTRv2, while improving 14.3 mAP.
+Our UVE accounts for only 15.3% of the total pipeline, highlighting the efficiency and practical value of our approach across various application scenarios.
+Appendix B More Qualitative Visualization
+We visualize map construction results of PriorDrive under various weather conditions on nuScenes in Figure 6.
+PriorDrive maintains stable and impressive results.
+Moreover, in Figure 7, Figure 8 and Figure 9, we show the result comparison between MapTRv2 and the proposed PriorDrive on the nuScenes.
+Finally, we also present the visualizations of Argoverse 2 in Figure 10 and OpenLane-V2 in Figure 11.
+Our results have more complete map representations, more accurate shape and point positions of map elements.
+Figure 6: Our method maintains stable and robust HD map construction quality in complex and variable driving scenarios on nuScenes.
+Figure 7: Qualitative result comparison under the weather condition of cloudy on nuScenes.
+Figure 8: Qualitative result comparison under the weather condition of rainy on nuScenes.
+Figure 9: Qualitative result comparison under the night condition on nuScenes.
+Figure 10: Qualitative results on Argoverse2 val set. Since we generate a 3D vectorized HD map, the predictions can be rendered precisely on the surrounding view images.
+Figure 11: Qualitative results on OpenLane-V2. The red arrows represent the predicted lane topology.